### v1

# Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Nonlinear Optical Metamaterial

Abstract
The systematic exploration of ABC type heterostructures reveals that nanoscale morphological modification markedly improves nonlinear optical properties to maximize the artificial bulk second-order susceptibility. These amorphous birefringent heterostructures are fabricated through cyclic plasma-enhanced atomic layer deposition of three oxides, effectively breaking centrosymmetry. We observe a dependence of optical nonlinearity on the thickness variation of three constituent materials: SiO2 (A), TiO2 (B), and Al2O3 (C), ranging from tens of nanometers to the atomic scale, and these materials exhibit second-order susceptibility at their interfaces. Our findings reveal that the enhancement of nonlinear optical properties is strongly correlated with a high density of layers and superior interface quality, where the interface second-order nonlinearity transitions to bulk-like second-harmonic generation. An effective bulk second-order susceptibility of $\chi_{zzz}=2.0\pm 0.2$ pm/V is achieved, comparable to typical values for conventional monocrystalline nonlinear materials.
second harmonic generation, surface second-order susceptibility, interface second-order susceptibility, nonlinear metamaterial, ABC heterostructure, nanolaminates, quaternary oxide, tricolour heterostructure
I Introduction
example2023The rise of atomic layer deposition (ALD) has revolutionized numerous technologies, including transistors ALD-tranzistor , memory devices ALD-memory , solar cells solarcell , catalysts catalyst , batteries batteries , conformal optical coatings AdrianaALD , and many others. However, its impact on the field of nonlinear optics remains largely untapped. Our research investigates the potential of ALD-deposited amorphous heterostructures, specifically ABC type nanolaminates consisting of three distinct dielectrics SiO2 (A), TiO2 (B), and Al2O3 (C), to create CMOS-compatible second-order nonlinear metamaterials. The core concept of metamaterials is to engineer materials with properties that are not found in nature by designing their structure at a scale smaller than the wavelength of light they interact with. Atomic layer deposition enables unparalleled precision in the fabrication of the structures, building them one atomic layer at a time with exceptional control and uniformity. Previous studies on layered composites with nonzero bulk second-order susceptibility assumed that only non-centrosymmetric constituents contribute to nonlinearity. However, studies neglected interface contributions to the nonlinear susceptibility of the heterostructure Boyd:94 . Our goal is to leverage the often overlooked interface, or surface, second-order nonlinearity heinz1991second ; shen1989optical by stacking multiple thin oxide layers in 2D heterostructures, thus significantly enhancing the effective nonlinear response ABC-Alloatti ; ABC-Clemmen . Impressively, these heterostructures exhibit a robust, bulk-like nonlinear response, even though all constituent oxides lack inherent bulk second-order nonlinearity and are amorphous, as shown in Fig. 1. This produces a highly distinctive material where second-order nonlinear processes arise not from bulk nonlinearity but from surface nonlinearity at the interfaces between centrosymmetric constituents. Previous studies on ternary oxide heterostructures (Al2O3/TiO2 and SiO2/HfO2) have demonstrated superior structural and optical properties of ALD-grown thin films, achieving precise control down to the atomic scale palabi ; PallabiHfO2 . These structures offer temperature and chemical stability, low-temperature growth requirements of $100\penalty 10000\ ^{\circ}\text{C}$, and conformal deposition on a wide range of nano- and micro-structured substrates, without the need to account for substrate compatibility or crystal growth orientation, unlike nonlinear III–V semiconductors such as GaAs GaAs . Moreover, they bypass the high-temperature manufacturing requirements of lithium niobate on insulator (LNOI), another widely used nonlinear platform LNOI . They also provide the flexibility to engineer the effective refractive index for specific applications by adjusting the ratio of constituent oxides palabi . ABC stacks exhibit outstanding optical transparency across the visible spectrum, surpassing the performance of typical III-V technologies. Moreover, the ALD deposition technique and the dielectrics used are standard in the semiconductor industry. Altogether, these advantages make ABC heterostructures a powerful platform for integrating photonic structures on a chip using only commonly available oxides and established semiconductor processes.
Figure 1: The visualization of the ABC type heterostructures with an optical axis perpendicular to the layers. Changing the density of layers influences nonlinear optical properties.
In bulk media with inversion symmetry, second-order nonlinear processes are forbidden under the electric dipole approximation, leading to zero bulk second-order susceptibility, $\chi_{\text{bulk}}^{(2)}=0$ Shen ; Boyd1992 , and no second harmonic enhancement is observed with varying thickness in pure samples of centrosymmetric materials A, B, or C. However, at the interface between two centrosymmetric media, the inversion symmetry, defined by the transformation $(x,y,z)\to(-x,-y,-z)$, is broken. This symmetry breaking results in a nonzero surface second-order susceptibility, $\chi_{\text{s}}^{(2)}\neq 0$, enabling electric-dipole-allowed second-harmonic generation (SHG) at surfaces and interfaces Shen1989 ; GuyotSionnest1987 . This principle is the key behind ABC type heterostructures, which consist of nanometer-thick amorphous (centrosymmetric) layers of three materials: A, B, and C. The inversion symmetry is broken locally at each AB, BC, and CA interface. In the case of AB stacks, the centrosymmetry is broken only locally, not globally. However, in ABC stacks, the inversion symmetry is broken both locally at the interfaces and globally, prohibiting total destructive interference of the second harmonic, resulting in a net $\chi_{\text{s}}^{\text{ABC}}$ for the ABC stack. The literature suggests that ABCD stacks, such as those incorporating SiO2/TiO2/Al2O3 with the addition of Hf2O2, may outperform ABC stacks. However, this potential enhancement remains experimentally unverified ABC-Alloatti ; ABC-Clemmen . Unfortunately, clear studies are lacking on which combination of inversion-symmetric materials offers the best nonlinear responses. A promising approach is to select materials with high dielectric contrast heinz1991second . However, experimental values for surface second-order susceptibility of materials and interfaces are scarce rodriguez2008calibration ; wang2010multipolar .
Surface nonlinearity arises from three main physical origins. First, the electric dipole contribution due to symmetry breaking at interfaces. Second, a field discontinuity at the interface, where the normal component of the electric field ($E_{z}$) changes rapidly, contributes to the surface nonlinearity via the electric quadrupole (nonlocal) term. Third, structural disparities between materials, such as liquid/solid interfaces, contribute to bulk-like surface nonlinearity even when dielectric constants are matched. The second and third contributions are known as electric quadrupole or nonlocal contributions to surface nonlinear susceptibility shen1989optical . Electric dipoles dominate second-harmonic generation, although small contributions from electric quadrupoles and magnetic dipoles are not ruled out multipolar1 ; multipolar2 .
ABC type heterostructures have birefringent properties hermans2019 . For monocrystals, birefringence can be explained by the anisotropic electrical properties of the molecules that make up the crystals. However, birefringence can also be caused by anisotropy if the material is arranged on an order of magnitude larger than the size of the molecules and this distance is still smaller than the wavelength of light. The ABC type heterostructure, composed of three materials arranged in thin layers, acts as an effective medium exhibiting form birefringence and functioning as a negative uniaxial crystal. At the same time, the entire structure is amorphous, and the optical axis is perpendicular to the layers born2013principles ; Formbiro .
In this work, we present experimental investigations of ABC heterostructures in the form of SiO2/TiO2/Al2O3, deposited via atomic layer deposition. We systematically vary the thickness of each constituent material to evaluate its impact on the efficiency of second-harmonic generation. We systematically explore the structural, linear, and nonlinear optical properties for each morphological composition.
II Deposition and Characterization
Atomic layer deposition
Plasma-enhanced atomic layer deposition (PEALD) was performed using a SILAYO-ICP330 system (Sentech Instruments GmbH) to fabricate SiO2/TiO2/Al2O3 heterostructures. The deposition utilized oxygen plasma as the co-reactant with the following precursors: bis(diethylamino)silane (BDEAS) for SiO2, tetrakis(dimethylamino)titanium (TDMAT) for TiO2, and trimethylaluminum (TMA) for Al2O3. The inductively coupled plasma (ICP) source was operated at $100\penalty 10000\ \text{W}$ for all processes. The reactor chamber was kept at a temperature of $100\penalty 10000\ ^{\circ}\text{C}$. A 15-minute stabilization period was implemented after sample loading prior to deposition to ensure thermal stability and uniformity. Nitrogen (N2) served as the precursor carrier and purge gas. The growth per cycle (GPC) is determined for single-layer depositions of approximately 20-50 nm. All deposition parameters are summarized in Table 1.
Table 1: PEALD process parameters for SiO2/TiO2/Al2O3 heterostructures.
Material SiO2 TiO2 Al2O3
Refractive index 1.45 2.29 1.63
Precursor BDEAS TDMAT TMA
Pulse (ms) 300 3130 80
Purge (ms) 5000 8000 2000
Plasma Gas O2 O2 O2
Pulse (ms) 3000 5000 3000
Purge (ms) 2000 5000 2000
ICP power (W) 100 100 100
O2 flow (sccm) 200 200 100
N2 flow (sccm) 30 160 80
[…]
Growth per cycle ($\mathrm{\SIUnitSymbolAngstrom}$/cycle) 1.2 0.7 1.5
sccm: standard cubic centimeter per minute.
X-ray reflectivity
[…]
Transmission electron microscopy
To examine the samples, high-resolution scanning transmission electron microscopy (HR-STEM) studies were carried out using a Themis Z(3.1) 80-300 TEM instrument (Thermo Fisher Scientific) with probe Cs-aberration correction, operating at 300 nm. The microscope is equipped with a Super-XTM (TFS) Energy Dispersive X-ray (EDX) detector and an a Gatan Image Filter (GIF) Continuum 1065ER used for Electron Energy Loss Spectroscopy (EELS). The samples were prepared by Focused Ion Milling (FIB) lift-out technique using a Helios G4 UC FIB-SEM device operated at 30, 16 and 5 kV.
Ellipsometry
The refractive indices were determined for each structure using an SE850 DUV variable angle spectroscopic ellipsometer (Sentech Instruments GmbH), and the data were evaluated in the wavelength range from 400nm400\ $\text${nm} to 1040nm1040\ $\text${nm} with a uniaxial Sellmeier dispersion model.
Nonlinear setup
The following optical setup was used to investigate the nonlinear optical response of the ABC layers. The femtosecond laser (PHAROS-SP, Light Conversion) emits pulses with full-width half-maxima (FWHM) of $177\text{\,}\mathrm{fs}$ with a central wavelength of 1032 nm and with an average power of up to $1.5\text{\,}\mathrm{W}$ at a repetition rate of $200\text{\,}\mathrm{kHz}$. Polarization-based measurements were performed by a femtosecond Satsuma laser (Amplitude) with full-width half maxima of $270\text{\,}\mathrm{fs}$, wavelength of 1031.4 nm and the repetition rate of $1\text{\,}\mathrm{MHz}$, Fig. 2. Power and polarization are controlled by a half-wave plate HW1 (WPH05M-1030, Thorlabs), polarizer P1 (VA5, Thorlabs), and half-wave plate HW2 (Thorlabs AHWP05M-950). The laser beam is focused on the sample by the plano-convex lens L1 (LA1172-B-ML, Thorlabs) with a focal length of 400 mm, producing a beam waist of $w_{0}=55\ \upmu\text{m}$. This leads to the confocal parameter of 18, that is, $2\timesRayleigh$ length. The loosely focused beam enables the front and back sides of the sample to remain in focus during the rotation process; the sample is 1 mm thick. A long pass filter LP (FELH0900, Thorlabs) is positioned behind the lens L1 to remove any parasitic light, especially any SHG created at other optical elements. The sample is mounted onto an automated rotation stage, and its position is precisely aligned so that the laser beam stays focused on the same position on the sample during the whole rotation. The beam spot must always lie on the y-axis of rotation. The sample is rotated from the angle of incidence $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $85\text{\,}\mathrm{\SIUnitSymbolDegree}$ in steps of $0.25\text{\,}\mathrm{\SIUnitSymbolDegree}$. This angle setting verifies proper sample alignment and improves the reliability of the evaluation process. To isolate the SH signal from the fundamental, a short-pass filter SP (FESH0800, Thorlabs) is used to remove the fundamental frequency. A Rochon prism analyzer P2 (RPM10, Thorlabs) is required to determine the direction of SHG polarization. A bandpass filter BP (Thorlabs FBH520-40) centered at 520 nm with FWHM 40 nm suppresses parasitic light and higher harmonics. The SHG signal is focused with a biconvex lens L2 (LA1608-A, Thorlabs) with a focal length of 100 mm on the camera (CS165MU1/M, Zelux). The camera calibration factor for calculating the SHG’s power is determined separately by using a Beta-Barium Borate (BBO) crystal as a reference sample. This samples produces a sufficiently strong SHG signal, which can be measured by a calibrated photodiode (S121C, Thorlabs) placed on the camera place. Subsequently, several optical density filters were introduced in the optical path to lower the power of the SH from the BBO crystal to the same order of magnitude as the one from the sample. Knowing the power output and optical density of the filters, we can correlate the power recorded by the camera with the measured signal. The signal was summed up in the area of images with recorded SHG, subsequently divided by the exposure time, and multiplied by the calibration factor of the camera to determine the optical power.
Figure 2: Graphical schema of the optical setup used for second harmonic measurements of the ABC type heterostructures. From the left: femtosecond laser, HW1: half-wave plate, P1: polarizer, HW2: half-wave plate, L1: plano-convex lens, LP: long pass filter, rotation stage with the sample, SP: short pass filter, L2: biconvex lens, P2: Rochon prism analyzer, BP: bandpass filter, camera.
Nonlinear characterization
The standard method for evaluating nonlinear second-order susceptibility $\chi^{\textup{(2)}}$ is the Maker fringe technique Maker . It uses the rotation of the sample while the SHG intensity is collected for each angle. However, it is limited to thicker crystals with a minimum size of tens of micrometers to a few millimeters. The limiting factor led to the development of a new method for measuring second-order susceptibility by Hermans et al. On-determination-SHG ; hermans2019 ; koskinen2018 . This method is well suited for evaluating the $\chi^{(2)}$ values of thin films with thicknesses even less than 100 nm where phase mismatch effects can be neglected. Like the Maker fringe technique, the sample is rotated during SHG measurements using a fixed polarization fundamental frequency, and the obtained SH curve is used to fit the nonlinear coefficients. This method does not consider individual separated layers A, B, and C and each interface as a source of second harmonic, but uses an effective medium approximation. Neglecting its birefringent properties, the ABC heterostructure is represented by an average refractive index of $(n_{0}+n_{e})/2$. The SHG is described in terms of the surface second-order nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ of the total heterostructure.
Any second-order nonlinear process is described by the second-order susceptibility tensor $\chi^{\textup{(2)}}$, with 27 elements. However, the number of nonzero and independent elements is much smaller due to symmetry properties. The ABC heterostructure belongs to the symmetry group $C_{\infty v}$, as do other amorphous surfaces or thin achiral films with in-plane isotropy. This means that the nonzero coefficients of the second-order susceptibilities are $\chi^{\textup{(2)}}_{{xxz}}=\chi^{\textup{(2)}}_{{yyz}}=\chi^{\textup{(2)}}_{{zxz}}=\chi^{\textup{(2)}}_{{yzy}}$, $\chi^{\textup{(2)}}_{{zxx}}=\chi^{\textup{(2)}}_{{zyy}}$ and $\chi^{\textup{(2)}}_{{zzz}}$, where the $z$ direction is perpendicular to the layers and the $x$ and $y$ directions are longitudinal to the layers. It is also assumed that these coefficients have only real values and that the measurement takes place outside the resonance region, i.e. absorption is not present. The general description of the electric field components of the SH coming from a surface with $C_{\infty v}$ symmetry is described as:
$E_{2\omega,\textup{p}}=fE_{\omega,\textup{p}}^{2}+gE_{\omega,\textup{s}}^{2},$ (1)
$E_{2\omega,\textup{s}}=hE_{\omega,\textup{p}}E_{\omega,\textup{s}},$ (2)
where p and s represent p- and s-polarized components and $E_{\omega}$ is the incoming electric field amplitude of the fundamental frequency. The coefficients $f$, $g$, and $h$ represent generalized layer properties such as nonlinear susceptibility tensor components, thickness, refractive index, and angle of incidence.
The measurement technique is based on the fact that there are two sources of SH signals, $\chi^{\textup{ABC}}_{\textup{s}}$ at the air-thin film interface at the frontside of the sample and $\chi^{\textup{glass}}_{\textup{s}}$ at the glass-air interface at the backside of the sample. Both SH signals interfere with each other, producing specific angle-dependent fringes for this method. Based on prior knowledge of the magnitude of $\chi^{\textup{glass}}_{\textup{s}}$, which is known from either the literature or from calibration measurements, the magnitude of the unknown nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ in the thin layer can be deduced. We can obtain the bulk second-order susceptibility as $\chi^{\textup{ABC}}_{\textup{b}}=\chi^{\textup{ABC}}_{\textup{s}}/d$, where $d$ is the total thickness of the heterostructure. Under these approximations made by Hermans et al., a monochromatic p-polarized plane wave of the fundamental frequency of the amplitude of the electric field $E_{\omega,\textup{p}}$ produces a transmitted p-polarized SH electric field $E_{2\omega,\textup{total}}$ as
$E_{2\omega,\textup{total}}$ $=E_{2\omega,\textup{front}}+E_{2\omega,\textup{back}}$ (3)
$=-j\frac{\omega}{2c}t^{2}_{\textup{air,ABC}}E^{2}_{\omega,\textup{in}}\Bigg[\frac{T_{\textup{ABC,glass}}T_{\textup{glass,air}}}{N_{\textup{ABC}}\cos(\Theta_{\textup{ABC}})}\chi^{\textup{ABC}}_{\textup{s,eff}}\exp\left(-j\frac{2\omega N_{\textup{glass}}\cos(\Theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)$
$\quad-\frac{t^{2}_{\textup{ABC,glass}}T_{\textup{glass,air}}}{N_{\textup{glass}}\cos(\Theta_{\textup{glass}})}\chi^{\textup{glass}}_{\textup{s,eff}}\exp\left(-j\frac{2\omega n_{\textup{glass}}\cos(\theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)\Bigg],$
where $\chi^{{i}}_{\textup{s,eff}}$ is the effective surface second-order susceptibility
$\chi_{\textup{s,eff}}^{i}$ $=\chi_{\textup{s},xxz}^{i}\sin(2\theta_{i})\cos(\Theta_{i})+\chi_{\textup{s},zxx}^{i}\sin(\Theta_{i})\cos^{2}(\theta_{i})+\chi_{\textup{s},zzz}^{i}\sin^{2}(\theta_{i})\sin(\Theta_{i})$ (4)
$\approx\left(\chi_{\textup{s},zxx}^{i}+2\chi_{\textup{s},xxz}^{i}\right)\sin(\Theta_{i})\cos^{2}(\theta_{i})+\chi_{\textup{s},zzz}^{i}\sin^{2}(\theta_{i})\sin(\Theta_{i}).$
The parameters for the fundamental frequency (FF) $\omega$ are written in lower case letters, and for the SH in upper case. The parameters $T_{i,j}$ and $t_{i,j}$ are the Fresnel transmission coefficients for p-polarized light that transmits from medium $i$ to medium $j$. The variables $N_{i}$ and $n_{i}$ are refractive indices medium $i$. The variables $\Theta_{i}$ and $\theta_{i}$ are angles of propagation in each medium $i$ with respect to the surface normal $z$. The speed of light is $c$ and the thickness of the substrate is $L_{\textup{glass}}$. The refractive indices for the glass substrate made of pure fused silica are $N_{\textup{glass}}=1.4615$ and $n_{\textup{glass}}=1.4500$. Approximation in Eq. 4 is valid only for materials with low dispersion with $\Theta_{i}\approx\theta_{i}$. Due to this approximation, the values $\chi_{\textup{s},zxx}^{\textup{ABC}}$ and $\chi_{\textup{s},xxz}^{\textup{ABC}}$ cannot be fitted independently, but only as a sum $A_{\textup{s},zx}^{\textup{ABC}}=\chi_{\textup{s},zxx}^{\textup{ABC}}+2\chi_{\textup{s},xxz}^{\textup{ABC}}$. For substrate nonlinearity we used literature values Kauranen_glass_calibration at wavelength 1064 nm $\chi_{\textup{s},zxx}^{\textup{glass}}=3.8\times 10^{-22}$ pm/V, $\chi_{\textup{s},xxz}^{\textup{glass}}=7.9\times 10^{-22}$ pm/V, $\chi_{\textup{s},zzz}^{\textup{glass}}=5.9\times 10^{-22}$ pm/V and we apply the Miller’s rule
$\begin{aligned} \frac{\chi^{(2)}(2\omega)}{\chi^{(1)}(2\omega)\cdot\left|\chi^{(1)}(\omega)\right|^{2}}=\text{const.}\end{aligned},$ (5)
with $\chi^{(1)}$ as the linear susceptibility of the material, to estimate the values at a wavelength of 1032 nm. However, since a femtosecond laser is used, a delay is introduced between the two pulses from the front and back sides of the sample. An SH pulse is generated on the front side of the sample. When it reaches the back side of the sample, another SH pulse is already generated because the pulse of the fundamental wavelength travels at different group velocity through the substrate compared to the SH pulse. This delay between the two SH pulses when they reach the backside of the sample is called the temporal walk-off time $t_{\text{walk-off}}$. This causes destructive interference to be incomplete, and thus, the depth of the interference fringes is decreased. When the temporal walk-off is not taken into account, the value of $\chi_{\textup{}}^{(2)}$ is overestimated. For our fused silica sample, 1 mm thick, the value was measured to be $t_{\text{walk-off}}=94$ fs. This value corresponds well with previous measurements in the literature and also with the calculation of the delay based on the group velocity of each pulse in the substrate. Therefore, the model was modified by Hermans et al. for the pulsed laser with the temporal walk-off effect as
$\begin{aligned} P_{2\omega}=K_{2}\int_{-\infty}^{+\infty}\left[E_{2\omega,\text{front}}\ \text{sech}^{2}\left(\frac{t}{\frac{\Delta t}{2\ln\left(1+\sqrt{2}\right)}}\right)+E_{2\omega,\text{back}}\ \text{sech}^{2}\left(\frac{t+t_{\text{walk-off}}}{\frac{\Delta t}{2\ln\left(1+\sqrt{2}\right)}}\right)\right]^{2}\,dt\end{aligned},$ (6)
where $K_{2}$ is the calibration constant, $\Delta t$ is the pulse duration at FWHM. However, the walk-off time is not constant; it changes as the angle of incidence increases, and the optical path through the substrate increases as well. Then, the relation is $t_{\text{walk-off}}={t_{\text{walk-off,0}}}/{\cos(\theta_{\text{glass}})}$, where $t_{\text{walk-off,0}}$ is walk-off time at the normal angle of incidence. The walk-off effect is also visualized in Fig. 3. By substituting Eq. 6 with Eq. 3, we can obtain the desired second-order susceptibility together with the other fitting parameters $\chi_{\textup{s},zzz}^{\textup{ABC}}$, $A_{\textup{s},zx}^{\textup{ABC}}$, $K_{2}$, and $L_{\textup{glass}}$. The thickness of the glass substrate must also be fitted for each measurement, as the position of the fringes in the second harmonic is highly sensitive to variations in substrate thickness. Dividing the surface susceptibility by the total thickness of the heterostructure $\chi_{\textup{bulk}}^{\textup{ABC}}=\chi_{\textup{s}}^{\textup{ABC}}/d$, the bulk second-order susceptibility is determined. This simplified version of the model neglects multiple reflections in the film and substrate while maintaining validity, as previously tested. Therefore, we keep the angle of incidence below $60^{\circ}$ for the analysis.
The primary source of uncertainty in the $\chi^{(2)}$ evaluation stems from the uncertainty in determining the angle of incidence, estimated as a systematic error of $\pm 1^{\circ}$. This leads to an uncertainty in the $\chi^{(2)}$ value that is substantially greater than the standard deviations derived from the nonlinear regression analysis. We estimate the uncertainty to be approximately 10% for the stronger SH signal samples and up to 20% for the weaker ones.
The method currently has two limitations. First, it assumes no phase mismatch in the second harmonic, rendering it applicable only to thin films. Second, the SH signal originates not only from the ABC stack but also from the interfaces between the first deposited layer of material A and the substrate, and between the last deposited layer of material C and air. The method fails to account for these additional sources of second harmonic generation, which can lead to an overestimation of the second-order nonlinear susceptibility of the ABC heterostructure, particularly in samples with a limited number of ABC interfaces where the C-air interface may become significant. To estimate this effect, single layers of each material were also investigated.
Figure 3: (a) Visualizing the angle-dependent measurement for the analysis of second harmonic generation from ABC heterostructures with the angle of incidence $\vartheta$ and the axis orientation. (b) Showing effects present in the experiment. A stronger SH signal comes from the ABC layers, and it interferes constructively and destructively with the weaker SH signal from the back side of the substrate. The temporal walk-off effect also influences the interference by making the local minima shallower.
III Results and discussion
To investigate the nonlinear interface response of ABC heterostructures, multiple samples with a constant total deposition thickness of approximately 60 nm were fabricated, while systematically varying the ABC period thickness $t_{\text{ABC}}$ from 60 nm down to the atomic scale of 0.3 nm. This approach enabled the study of nonlinear responses by altering the density of interfaces per unit thickness and exploring different morphologies across the samples, with the goal of understanding how these variations influence the properties of heterostructures. Each oxide material is represented approximately equally in the period $t_{\text{ABC}}$. The ABC heterostructures consist of around 300–380 atomic layers for a 60 nm thick period and only about two atomic layers for an ABC period thickness of 0.3 nm, assuming a Si–O bond length of 0.15 nm, a Ti–O bond length of 0.20 nm, and an Al–O bond length of 0.16 nm Si-O ; Ti-O ; Al-O .
Scanning transmission electron microscope (STEM) images were collected using a high-angle annular dark-field (HAADF) and annular dark field (ADF) detectors to reveal the structural characteristics of the SiO2/TiO2/Al2O3 nanolaminates on a silicon wafer substrate in selected samples S3 and S2. The ADF images are shown in the Fig. 4. Sample S3 exhibits a clear ABC period thickness $t_{\text{ABC}}$ of 1.5 nm, with clearly visible distinct layers. On the other hand, sample S2, with a reduced $t_{\text{ABC}}$ of 0.7 nm, shows no recognizable layer separation, indicating a loss of structural distinction on this scale.
Figure 4: ABF STEM images showing the SiO2/TiO2/Al2O3 nanolaminates on the silicon wafer substrate of: (left) sample S3 with the ABC period thickness $t_{\textup{ABC}}$ of 1.5 nm; TiO2 layers appear darker; (right) sample S2 with $t_{\textup{ABC}}$ of 0.7 nm. The layers are no longer visible. The dark area on the top of both nanolaminates is the Pt/C protection layer used for TEM specimen preparation.
To verify the precise thickness of the layers in each composition, X-ray reflectivity (XRR) was employed. The measurement spectra of the heterostructures are presented in Fig. 5. The Bragg peaks are clearly visible for all samples except for the heterostructure with a period of 0.7 nm (sample S2). As the period of the heterostructure decreases, the first Bragg peak shifts to the higher grazing angle. The XRR analysis shows the existence of separated nanolaminates and the excellent agreement with the targeted ABC period thickness. The analysis also measured the total deposition thickness and provided the average interface width, incorporating both roughness and interlayer diffusion, see Table 2. Sample S2 has an interface width of 1.2 nm, greater than other samples, with a period thickness of $t_{\text{ABC}}=0.7$ nm, suggesting notable layer intermixing and a lack of clearly separated layers.
Figure 5: Results of the X-ray reflectivity measurements proving the existence of the layer separation and verifying the period thickness $t_{\text{ABC}}$. Sample ID from the top: S13, S11, S6, S3, S2.
Table 2: Results of the X-ray reflectivity measurements for selected SiO2/TiO2/Al2O3 heterostructures.
Sample ID Total thickness $d$ (nm) Period thickness $t_{\textup{ABC}}$ (nm) Layer thickness SiO2 (nm) Layer thickness TiO2 (nm) Layer thickness Al2O3 (nm) Interface width (nm)
S2 57.3 0.7 0.4 0.16 0.15 1.2
S3 59.7 1.5 0.4 0.4 0.7 0.4
S6 62.6 3.1 1.4 1.0 0.8 0.3
S11 59.9 6.0 1.5 1.7 2.8 0.4
S13 59.6 11.9 3.4 2.8 5.7 0.7
Table 3: Sample parameters for SiO2/TiO2/Al2O3 heterostructures and their optical properties.
Sample ID Deposition cycles SiO2/TiO2/Al2O3 Super cycles $M$ Layer thickness (nm) SiO2/TiO2/Al2O3 Period thickness $t_{\textup{ABC}}$ (nm) Total thickness $d$ (nm) $n_{\text{o}}$ (1032nm)(1032\ $\text${nm}) $n_{\text{e}}$ (1032nm)(1032\ $\text${nm}) $N_{\text{o}}$ (516nm)(516\ $\text${nm}) $N_{\text{e}}$ (516nm)(516\ $\text${nm}) $A_{\text{zx}}$ (pm/V) $\chi_{\text{zzz}}$ (pm/V)
[…]
S4 4:6:4 120 0.48/0.42/0.6 1.5 174 1.79 1.75 1.84 1.80 0.68 2.1
S5 4:6:4 160 0.48/0.42/0.6 1.5 233 1.78 1.76 1.84 1.80 0.77 2.5
[…]
S7 8:12:8 31 0.96/0.84/1.2 3.0 93 1.79 1.69 1.85 1.73 0.52 1.6
S8 8:12:8 39 0.96/0.84/1.2 3.0 119 1.79 1.69 1.85 1.73 0.55 1.6
S9 8:12:8 40 0.96/0.84/1.2 3.0 116 1.78 1.70 1.84 1.74 0.43 1.4
S10 8:12:8 70 0.96/0.84/1.2 3.0 215 1.78 1.71 1.84 1.75 0.69 1.9
[…]
S12 16:24:16 20 1.92/1.68/2.4 6.0 114 1.79 1.64 1.84 1.67 0.070 0.96
[…]
*: No birefringence was measured, $\dagger$: The effective medium approximation is no longer valid in the wavelength range from 400 nm to 1040 nm based on the ellipsometric evaluation Layer and period thickness $t_{\textup{ABC}}$ is calculated based on the growth rate per cycle, and the total thickness $d$ is based on ellipsometric evaluation. The uncertainty for the refractive index is $\pm 0.01$. Samples close to 60 nm of the total thickness are highlighted in gray.
The dispersion curve of the ordinary and extraordinary refractive index is presented in Fig. 6. The magnitude of birefringence, $\Delta n=n_{\text{e}}-n_{\text{o}}$, depends on the thickness of the ABC period, and completely disappears in samples with nonexistent layer separation. Optical properties are summarized in Table 3.
Figure 6: Ordinary and extraordinary refractive indices SiO2/TiO2/Al2O3 nanolaminates of the sample S3 show the presence of the form birefringence.
Polarization measurements were carried out to verify the symmetry properties of the ABC stacks. Figure 7 shows s- and p-polarization of the SH signal from sample S3 as a function of the polarization angle of the fundamental field $\varphi$ with the angle of incidence $\vartheta=63^{\circ}$. S-polarization is represented by $\varphi=0^{\circ}$ and p-polarization by $\varphi=90^{\circ}$. The polarization measurements correspond with the symmetry group $C_{\infty v}$. The maximum SH signal is observed when both the fundamental and the second harmonic are p-polarized.
Figure 7: Normalized SHG power for s- and p-polarizations as a function of the polarization angle $\varphi$ of the fundamental corresponding with the symmetry group $C_{\infty v}$, measured for the sample S3.
Based on the experimental results for the p-polarized fundamental and the second harmonic in Fig. 8(a), the intensity of the second harmonic increases with decreasing ABC period thickness from 60 nm until 1.5 nm, where a maximum is observed. For periods thinner than 1.5 nm, the second harmonic signal decreases abruptly. As each layer is approximately 0.25 nm thick, intermixing at the interfaces likely degrades layer quality, and incomplete layer closure further reduces the SH signal. This hypothesis is supported by STEM imaging and XRR analysis of sample S2, which show no clear layer separation for an ABC period of $t_{\textup{ABC}}=0.7$ nm. The nonlinear model for the determination of $\chi^{\textup{(2)}}$ is fitted to all SHG data across varying ABC periods, with selected curves shown in Fig. 8(b).
Figure 8: (a) Angle-dependent measurements of the second harmonic frequency from SiO2/TiO2/Al2O3 heterostructures deposited on 1 mm thick fused silica substrate. The total deposited thickness is approximately 60 nm across all samples; only the morphology, i.e., the ABC period thickness $t_{\textup{ABC}}$, varies from sample to sample. Both FF and SH are p-polarized with average power $P=0.5$ W. (b) Fitting of SHG data from selected samples to analyze the second-order susceptibility of the ABC heterostructures.
The same trend is also reported for power dependency measurements. The samples are positioned at the angle of incidence when SHG maxima are observed for p-polarized FF and SHG. In addition, an uncoated substrate of fused silica with natural surface second-order nonlinearity is also included to compare it with the SH intensity across the ABC samples. Verification of second-harmonic generation is achieved by observing its quadratic power dependency, and excellent agreement is found by achieving a slope line of $2.026(9)$ for sample S3, Fig. 9. The second harmonic exhibits a quadratic relationship with the ABC super cycles $M$, similar to the classical bulk crystals, Fig. 10. For each ABC period, second-order susceptibility is determined with the results shown in Fig. 11(a). The trend suggests an enhancement of both $\chi_{zzz}$ and $A_{zx}$ with decreasing ABC period thickness and achieving $\chi_{zzz}=2.0\pm 0.2$ pm/V for $t_{\textup{ABC}}=1.5$ nm. For shorter periods, the values of $\chi_{zzz}$ and $A_{zx}$ decrease rapidly. Several thicknesses were analyzed to clarify the properties at different deposited total thicknesses $d$. The analysis shows that the nonlinear susceptibility remains constant for total thicknesses from 60 nm to 233 nm for a period of 1.5 nm. A similar trend is also observed for periods of 3 nm and 6 nm.
Figure 9: Second harmonic power dependence on fundamental power in SiO2/TiO2/Al2O3 heterostructures, measured at the angle of incidence of maximum second harmonic intensity for both p-polarized fundamental and second harmonic. Quadratic power dependence of the second harmonic for sample S3 with period $t_{\textup{ABC}}=1.5$ nm, fitted to the data.
Figure 10: Second harmonic power scales with quadratic dependence on the film thickness and with the number of ABC super cycles $\text{ABC}]\times M$. Both fundamental and SH are p-polarized with average fundamental power $P=0.5$ W. ABC period thickness is $t_{\textup{ABC}}=3.0$ nm for all samples. Measured at the angle of incidence $\vartheta$ where the global maxima occurs.
Figure 11: (a) Analysis of second-order susceptibility $\chi_{zzz}$ and $A_{zx}$ as a function of ABC period thickness. (b) Investigating second-order susceptibility $\chi_{zzz}$ across variations in ABC period and total film thickness.
Our best ABC compositions are compared in Table 4 with other commonly used nonlinear crystals, as well as with previously reported results for ABC type heterostructures in the literature. We achieve higher nonlinearity than previously reported values, and we attribute this improvement to better-optimized PEALD processes. For example, Alloatti et al. ABC-Alloatti report difficulties with the HfO2 process. Moreover, we achieved much thinner and still enclosed layers compared to the previous reports. ABC period thicknesses $t_{\textup{ABC}}$ were 2.1 nm and 2.7 nm, compared to our best sample with 1.5 nm. Material selection may be partially responsible for the higher observed nonlinearity caused by the dielectric contrast between the layers.
Table 4: Comparing ABC-type heterostructures with other nonlinear monocrystals
Structure Material ABC material $n$ (-) $n_{\textup{o}}$$n_{\textup{e}}$ Point Group $\lambda$ (nm) $t_{\textup{ABC}}$ (nm) $d$ (pm/V) Ref.
Monocrystal SiO2 (Quartz) - 1.53/1.54 $\textit{D}_{3}$ 1064 - $d_{\text{11}}$ = 0.3 lithiumbiobate ; quartz-n
Monocrystal BaB2O4 (BBO) - 1.67/1.54 $\textit{C}_{3v}$ 1064 - $d_{\text{22}}$ = 2.3 lithiumbiobate ; BBO-n
Monocrystal LiNbO3 (LN) - 2.23/2.16 $\textit{C}_{3v}$ 1064 - $d_{\text{33}}$ = -27 lithiumbiobate ; LN-n
ABC Amorphous HfO2/TiO2/Al2O3 2.0/2.1/1.6 - $C_{\infty v}$ 800 2.7 $d_{\text{33}}$ = 0.4 On-determination-SHG
ABC Amorphous In2O3/TiO2/Al2O3 2.2/2.1/1.6 - $C_{\infty v}$ 800 2.1 $d_{\text{33}}$ = 0.6 On-determination-SHG
ABC Amorphous (S3) SiO2/TiO2/Al2O3 1.45/2.33/1.62 1.77/1.75 $C_{\infty v}$ 1032 1.5 $d_{\text{33}}$ = 1.0 This work
ABC Amorphous (S6) SiO2/TiO2/Al2O3 1.45/2.33/1.62 1.79/1.71 $C_{\infty v}$ 1032 3.0 $d_{\text{33}}$ = 0.9 This work
By convention, $\chi^{(2)}=2d$
To further clarify the origin of the effective bulk nonlinearity, two identical ABC samples S3 were placed on top of each other. First, two samples with stacking sequence I) CBA…CBA-Silica/CBA…CBA-Silica and sequence II) CBA…CBA-Silica/Silica-ABC…ABC were investigated, Fig. 12. Both stacking sequences produce almost the same maximum SHG intensity. However, the number of interference fringes is increased for sequence b), and the depth of fringes is reduced. The temporal walk-off effect is more significant because the optical thickness of the substrates between the heterostructures is twice as large. However, when the stacking sequence I) Silica-ABC…ABC/CBA…CBA-Silica is applied, the SHG almost vanishes. This result suggests that the centrosymmetry is restored. However, SH waves must be generated in each ABC and CBA layer; therefore, they are generated in each stack with an opposite phase and destructively interfere. This leads to the conclusion that the two susceptibilities of the ABC and CBA layers must have an opposite sign, for example, $|\chi_{\textup{s}}^{\textup{ABC}}|=|-\chi_{\textup{s}}^{\textup{CBA}}|$. The polarity flipping of the second-order susceptibility by reversing the growing order was also demonstrated in semiconductor heterostructures based on quantum-engineered intersubband transitions, where the polarity is flipped by spatial inversion in the growth sequence chi2flipping , similar to our case. However, the ABC and CBA sequences can differ under real conditions. For example, material A can grow on material B at a different rate and quality than material B on A. It is yet to be determined whether all growth permutations ABC, ACB, BAC, BCA, CAB, and CBA result in equivalent stack characteristics.
Figure 12: (a) Showing each stacking sequence for the two identical samples with total deposited ABC stack thickness of 60 nm and thickness of the fused silica substrate is 1 mm. (b) Showing angle-dependent measurements for three distinct stacking sequences of the ABC sample, both for p-polarized fundamental and second harmonic. Depending on the stacking sequence configuration, the SH response varies significantly.
IV Conclusion
We demonstrated an enhancement of the effective bulk second-order susceptibility in ABC type heterostructures by optimizing the thickness of the ABC period, achieving an optimal value of $t_{\text{ABC}}=1.5$ nm and a dominant tensor component of $\chi_{zzz}=2.0\pm 0.2$ pm/V, based on the second-order nonlinear response at the interfaces. The observed non-linearity enhancement stems from the higher density of interfaces, symmetry breaking in the ABC stacks, and dielectric contrast between the materials. The nonlinearity could be further enhanced if the layers remain distinct and non-intermixed at sub-0.5 nm thicknesses, for example, by minimizing surface and interface roughness through deposition on ultra-smooth substrates. Experimental results indicate that the optimal strategy for maximizing second-order susceptibility is to incorporate the greatest number of ABC layers per unit thickness, while ensuring physical separation to maintain structural integrity. However, the layer thickness is ultimately constrained by the limit of a single atomic layer. Therefore, future efforts should prioritize identifying a more suitable material composition, which remains challenging and elusive. Additionally, ABC nanolaminates have potential for near-ultraviolet applications, leveraging their low absorption edge when constituent materials are carefully selected. Polarity flipping, combined with form birefringence, enables phase matching in bulk form-birefringent metamaterials formBirometamaterial . We envision that ABC type heterostructures will open new avenues for nonlinear optics by integration into nanophotonic waveguides to efficiently access the dominant tensor component $\chi_{zzz}$ ABC-waveguide , metasurfaces and other photonic platforms ABC-overcoating ; ABC-metasrufaceplasmonic .
Funding
This research received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the International Research Training Group (IRTG) 2675 “Meta-ACTIVE” (project number 437527638) and the Collaborative Research Center (CRC) 1375 “NOA.” Additional support was provided by the Fraunhofer Internal Programs under Grant No. SME 431 40-04871. We also gratefully acknowledge funding from the Federal Ministry of Research, Technology and Space under Project ID 13N16897 (Gradient). Authors also thank the European Regional Development Fund and the State of Brandenburg for the Themis Z microscope (part of Potsdam Imaging and Spectral Analysis (PISA) facility).
Disclosures
The authors declare no conflicts of interest.
Data Availability
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Appendix A Appendix
The following method On-determination-SHG was used to experimentally obtain the temporal walk-off time for our laser and 1 mm thick fused silica substrates. The femtosecond laser with p-polarization of the fundamental at a power of 1 W was used to measure the angle-dependent p-polarized SH signal from the front and back surfaces with known nonlinear susceptibility $\chi_{\textup{s}}^{\textup{glass}}$ on each side. Then, the equation is as follows:
$E_{2\omega,\textup{total}}$ $=E_{2\omega,\textup{front}}+E_{2\omega,\textup{back}}$ (7)
$=-j\frac{\omega}{2cN_{\textup{glass}}\cos(\Theta_{\textup{glass}})}t^{2}_{\textup{air,glass}}T_{\textup{glass,air}}E^{2}_{\omega,\textup{in}}$
$\ \ \ \ \ \left(\chi_{\textup{s},xxz}^{\textup{glass}}\sin(2\theta_{\textup{glass}})\cos(\Theta_{\textup{glass}})+\chi_{\textup{s},zxx}^{\textup{glass}}\sin(\Theta_{\textup{glass}})\cos^{2}(\theta_{\textup{glass}})+\chi_{\textup{s},zzz}^{\textup{glass}}\sin^{2}(\theta_{\textup{glass}})\sin(\Theta_{\textup{glass}})\right)$
$\ \ \ \ \Bigg[\exp\left(-j\frac{2\omega N_{\textup{glass}}\cos(\Theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)-\exp\left(-j\frac{2\omega n_{\textup{glass}}\cos(\theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)\Bigg].$
After substituting into the Eq. 6, we fit the parameters $K_{2}$, $L_{\textup{glass}}$ and $t_{\text{walk-off}}$. The fitting is restricted to the angle of incidence of 60° to limit the effect of multiple reflections in the sample. The obtained value for the temporal walk-off time is $t_{\text{walk-off}}=94$ fs, see Fig. 13.
Figure 13: The result of the fitting for the temporal walk-off time $t_{\text{walk-off}}$ for 1 mm fused silica substrate.
Other parameters for the evaluation of the second-order susceptibility are shown in the Table 5.
Table 5: Additional fitted parameters of SiO2/TiO2/Al2O3 heterostructures for the evaluation of the second-order susceptibility
Sample ID Calibration constant $K_{2}$ (-) $\times 10^{3}$ $L_{\text{glass}}$ ($m$)
S1 26 1.039
S2 88 1.010
S3 92 1.019
S4 83 0.999
S5 60 0.985
S6 95 1.043
S7 104 1.048
S8 101 1.027
S9 100 0.985
S10 70 1.025
S11 103 1.017
S12 88 0.999
S13 108 1.026
S14 90 1.017

### v4

@@
-Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Nonlinear Optical Metamaterial
-Abstract
-The systematic exploration of ABC type heterostructures reveals that nanoscale morphological modification markedly improves nonlinear optical properties to maximize the artificial bulk second-order susceptibility.
-These amorphous birefringent heterostructures are fabricated through cyclic plasma-enhanced atomic layer deposition of three oxides, effectively breaking centrosymmetry.
-We observe a dependence of optical nonlinearity on the thickness variation of three constituent materials: SiO2 (A), TiO2 (B), and Al2O3 (C), ranging from tens of nanometers to the atomic scale, and these materials exhibit second-order susceptibility at their interfaces.
+Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Second-Order Nonlinear Optical Metamaterial
+Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Second-Order Nonlinear Optical Metamaterial
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+journal: opticajournalarticletype: Research Article
+{abstract*}
+Systematic exploration of amorphous ABC heterostructures revealed that nanoscale morphological modifications markedly improved their artificial bulk second-order susceptibility.
+These amorphous birefringent heterostructures were fabricated through plasma-enhanced atomic layer deposition of three oxides, effectively breaking the centrosymmetry on the nanoscale.
+We observe a dependence of the optical nonlinearity on the thickness variation of three constituent materials, SiO2, TiO2, and Al2O3, ranging from tens of nanometers to the atomic scale, and these thin films exhibit second-order susceptibility at their interfaces.
 Our findings reveal that the enhancement of nonlinear optical properties is strongly correlated with a high density of layers and superior interface quality, where the interface second-order nonlinearity transitions to bulk-like second-harmonic generation.
-An effective bulk second-order susceptibility of $\chi_{zzz}=2.0\pm 0.2$ pm/V is achieved, comparable to typical values for conventional monocrystalline nonlinear materials.
-second harmonic generation, surface second-order susceptibility, interface second-order susceptibility, nonlinear metamaterial, ABC heterostructure, nanolaminates, quaternary oxide, tricolour heterostructure
-I Introduction
-example2023The rise of atomic layer deposition (ALD) has revolutionized numerous technologies, including transistors ALD-tranzistor , memory devices ALD-memory , solar cells solarcell , catalysts catalyst , batteries batteries , conformal optical coatings AdrianaALD , and many others.
-However, its impact on the field of nonlinear optics remains largely untapped.
-Our research investigates the potential of ALD-deposited amorphous heterostructures, specifically ABC type nanolaminates consisting of three distinct dielectrics SiO2 (A), TiO2 (B), and Al2O3 (C), to create CMOS-compatible second-order nonlinear metamaterials.
+An effective bulk second-order susceptibility of $\chi_{zzz}=2.0\pm 0.2$ pm/V at the wavelength of 1032 nm is achieved, comparable to typical values for conventional monocrystalline nonlinear materials.
+1 Introduction
+The rise of atomic layer deposition (ALD) has revolutionized numerous technologies, including transistors [44], memory devices [25], solar cells [40], catalysts [26], batteries [22], and conformal optical coatings [28].
+However, its impact on the field of nonlinear optics remains largely untapped [2, 7].
+Here, we investigate the potential of ALD-deposited amorphous heterostructures, specifically ABC-type nanolaminates consisting of three distinct dielectrics SiO2 (A), TiO2 (B), and Al2O3 (C) as shown in Fig.
+1, to create CMOS-compatible second-order nonlinear metamaterials.
 The core concept of metamaterials is to engineer materials with properties that are not found in nature by designing their structure at a scale smaller than the wavelength of light they interact with.
-Atomic layer deposition enables unparalleled precision in the fabrication of the structures, building them one atomic layer at a time with exceptional control and uniformity.
+Atomic layer deposition enables unparalleled precision in the fabrication of the thin films, building one atomic layer at a time with exceptional control and uniformity.
 Previous studies on layered composites with nonzero bulk second-order susceptibility assumed that only non-centrosymmetric constituents contribute to nonlinearity.
-However, studies neglected interface contributions to the nonlinear susceptibility of the heterostructure Boyd:94 .
-Our goal is to leverage the often overlooked interface, or surface, second-order nonlinearity heinz1991second ; shen1989optical by stacking multiple thin oxide layers in 2D heterostructures, thus significantly enhancing the effective nonlinear response ABC-Alloatti ; ABC-Clemmen .
-Impressively, these heterostructures exhibit a robust, bulk-like nonlinear response, even though all constituent oxides lack inherent bulk second-order nonlinearity and are amorphous, as shown in Fig.
-1.
+However, they neglected interface contributions to the nonlinear susceptibility of the heterostructure [6].
+Our goal is to leverage the often overlooked interface, or surface, second-order nonlinearity [13, 36] by stacking three thin oxide layers in 2D heterostructures, thus significantly enhancing the effective nonlinear response by breaking the centrosymmetry as discussed by L.
+Alloatti et al.
+and S.
+Clemmen at al.
+[2, 7].
+Impressively, these heterostructures exhibit a robust, bulk-like nonlinear response, even though all the constituent oxides lack inherent bulk second-order nonlinearity and are amorphous.
 This produces a highly distinctive material where second-order nonlinear processes arise not from bulk nonlinearity but from surface nonlinearity at the interfaces between centrosymmetric constituents.
-Previous studies on ternary oxide heterostructures (Al2O3/TiO2 and SiO2/HfO2) have demonstrated superior structural and optical properties of ALD-grown thin films, achieving precise control down to the atomic scale palabi ; PallabiHfO2 .
-These structures offer temperature and chemical stability, low-temperature growth requirements of $100\penalty 10000\ ^{\circ}\text{C}$, and conformal deposition on a wide range of nano- and micro-structured substrates, without the need to account for substrate compatibility or crystal growth orientation, unlike nonlinear III–V semiconductors such as GaAs GaAs .
-Moreover, they bypass the high-temperature manufacturing requirements of lithium niobate on insulator (LNOI), another widely used nonlinear platform LNOI .
-They also provide the flexibility to engineer the effective refractive index for specific applications by adjusting the ratio of constituent oxides palabi .
-ABC stacks exhibit outstanding optical transparency across the visible spectrum, surpassing the performance of typical III-V technologies.
-Moreover, the ALD deposition technique and the dielectrics used are standard in the semiconductor industry.
-Altogether, these advantages make ABC heterostructures a powerful platform for integrating photonic structures on a chip using only commonly available oxides and established semiconductor processes.
-Figure 1: The visualization of the ABC type heterostructures with an optical axis perpendicular to the layers. Changing the density of layers influences nonlinear optical properties.
-In bulk media with inversion symmetry, second-order nonlinear processes are forbidden under the electric dipole approximation, leading to zero bulk second-order susceptibility, $\chi_{\text{bulk}}^{(2)}=0$ Shen ; Boyd1992 , and no second harmonic enhancement is observed with varying thickness in pure samples of centrosymmetric materials A, B, or C.
+Previous studies on ternary oxide heterostructures (Al2O3/TiO2 and SiO2/HfO2) have demonstrated the superior structural and optical properties of ALD-grown thin films [27, 1].
+These structures offer temperature and chemical stability, low-temperature growth requirements of $100\penalty 10000\ ^{\circ}\text{C}$, and conformal deposition on a wide range of nano- and micro-structured substrates, without the need to account for substrate compatibility or crystal growth orientation, unlike nonlinear III–V semiconductors such as GaAs [19].
+Moreover, they bypass the high-temperature manufacturing requirements of lithium niobate on insulator (LNOI), another widely used nonlinear platform [29].
+They also provide the flexibility to engineer the effective refractive index for specific applications by adjusting the ratio of the constituent oxides [27].
+Oxide heterostructures exhibit outstanding optical transparency across the visible spectrum to UV spectral range, surpassing the performance of typical III-V technologies.
+Moreover, the ALD deposition technique and dielectrics used are standard in the semiconductor industry.
+Altogether, these advantages make ABC heterostructures a powerful platform for integrating photonic structures on a chip using only the commonly available oxides and established semiconductor processes.
+Figure 1: The visualization of the ABC-type heterostructures with an optical axis perpendicular to the layers. Changing the density of layers influences their nonlinear optical properties.
+In bulk media with inversion symmetry, second-order nonlinear processes are forbidden under the electric dipole approximation, leading to zero bulk second-order susceptibility, $\chi_{\text{bulk}}^{(2)}=0$ [37, 5], and no second harmonic enhancement is observed with varying thickness in pure samples of centrosymmetric materials A, B, or C.
 However, at the interface between two centrosymmetric media, the inversion symmetry, defined by the transformation $(x,y,z)\to(-x,-y,-z)$, is broken.
-This symmetry breaking results in a nonzero surface second-order susceptibility, $\chi_{\text{s}}^{(2)}\neq 0$, enabling electric-dipole-allowed second-harmonic generation (SHG) at surfaces and interfaces Shen1989 ; GuyotSionnest1987 .
-This principle is the key behind ABC type heterostructures, which consist of nanometer-thick amorphous (centrosymmetric) layers of three materials: A, B, and C.
+This symmetry breaking results in a nonzero surface second-order susceptibility, $\chi_{\text{s}}^{(2)}\neq 0$, enabling electric-dipole-allowed second-harmonic generation (SHG) at surfaces and interfaces [35, 11].
+ABC heterostructures make use of this principle by stacking nanometer-thick amorphous layers of three centrosymmetric materials A, B, and C.
 The inversion symmetry is broken locally at each AB, BC, and CA interface.
-In the case of AB stacks, the centrosymmetry is broken only locally, not globally.
+In the case of AB stacks, the centrosymmetry is broken only locally and not globally.
 However, in ABC stacks, the inversion symmetry is broken both locally at the interfaces and globally, prohibiting total destructive interference of the second harmonic, resulting in a net $\chi_{\text{s}}^{\text{ABC}}$ for the ABC stack.
-The literature suggests that ABCD stacks, such as those incorporating SiO2/TiO2/Al2O3 with the addition of Hf2O2, may outperform ABC stacks.
-However, this potential enhancement remains experimentally unverified ABC-Alloatti ; ABC-Clemmen .
+The literature suggests that ABCD stacks, such as those incorporating SiO2/TiO2/Al2O3 with the addition of HfO2, may outperform ABC stacks.
+However, this potential enhancement remains experimentally unverified [2, 7].
 Unfortunately, clear studies are lacking on which combination of inversion-symmetric materials offers the best nonlinear responses.
-A promising approach is to select materials with high dielectric contrast heinz1991second .
-However, experimental values for surface second-order susceptibility of materials and interfaces are scarce rodriguez2008calibration ; wang2010multipolar .
+A promising approach is to select materials with high dielectric contrast [13].
+However, experimental values for surface second-order susceptibility of materials and interfaces are scarce [32, 41].
 Surface nonlinearity arises from three main physical origins.
@@
 Third, structural disparities between materials, such as liquid/solid interfaces, contribute to bulk-like surface nonlinearity even when dielectric constants are matched.
-The second and third contributions are known as electric quadrupole or nonlocal contributions to surface nonlinear susceptibility shen1989optical .
-Electric dipoles dominate second-harmonic generation, although small contributions from electric quadrupoles and magnetic dipoles are not ruled out multipolar1 ; multipolar2 .
-ABC type heterostructures have birefringent properties hermans2019 .
-For monocrystals, birefringence can be explained by the anisotropic electrical properties of the molecules that make up the crystals.
-However, birefringence can also be caused by anisotropy if the material is arranged on an order of magnitude larger than the size of the molecules and this distance is still smaller than the wavelength of light.
-The ABC type heterostructure, composed of three materials arranged in thin layers, acts as an effective medium exhibiting form birefringence and functioning as a negative uniaxial crystal.
-At the same time, the entire structure is amorphous, and the optical axis is perpendicular to the layers born2013principles ; Formbiro .
-In this work, we present experimental investigations of ABC heterostructures in the form of SiO2/TiO2/Al2O3, deposited via atomic layer deposition.
+The second and third contributions are known as electric quadrupole or nonlocal contributions to surface nonlinear susceptibility [36, 42, 18].
+In this work, we present experimental investigations of ABC heterostructures in the form of SiO2/TiO2/Al2O3, deposited by atomic layer deposition.
 We systematically vary the thickness of each constituent material to evaluate its impact on the efficiency of second-harmonic generation.
-We systematically explore the structural, linear, and nonlinear optical properties for each morphological composition.
-II Deposition and Characterization
+The structural properties are studied by X-ray reflectivity, scanning transmission electron microscope, and time-of-flight secondary ion mass spectrometry.
+The film thickness and linear optical properties are determined by spectroscopic ellipsometry.
+Detailed study of second harmonic generation is carried out to evaluate the second-order nonlinear susceptibility for each morphological composition.
+2 Deposition and Characterization
 Atomic layer deposition
 Plasma-enhanced atomic layer deposition (PEALD) was performed using a SILAYO-ICP330 system (Sentech Instruments GmbH) to fabricate SiO2/TiO2/Al2O3 heterostructures.
-The deposition utilized oxygen plasma as the co-reactant with the following precursors: bis(diethylamino)silane (BDEAS) for SiO2, tetrakis(dimethylamino)titanium (TDMAT) for TiO2, and trimethylaluminum (TMA) for Al2O3.
+The deposition utilized oxygen plasma as a co-reactant with the following precursors: bis(diethylamino)silane (BDEAS) for SiO2, tetrakis(dimethylamino)titanium (TDMAT) for TiO2, and trimethylaluminum (TMA) for Al2O3.
 The inductively coupled plasma (ICP) source was operated at $100\penalty 10000\ \text{W}$ for all processes.
-The reactor chamber was kept at a temperature of $100\penalty 10000\ ^{\circ}\text{C}$.
+The reactor chamber was maintained at a temperature of $100\penalty 10000\ ^{\circ}\text{C}$.
 A 15-minute stabilization period was implemented after sample loading prior to deposition to ensure thermal stability and uniformity.
@@
 The growth per cycle (GPC) is determined for single-layer depositions of approximately 20-50 nm.
-All deposition parameters are summarized in Table 1.
+The deposition parameters are presented in Table 1.
 Table 1: PEALD process parameters for SiO2/TiO2/Al2O3 heterostructures.
 Material SiO2 TiO2 Al2O3
-Refractive index 1.45 2.29 1.63
+Refractive index (1032 nm) 1.45 2.29 1.63
 Precursor BDEAS TDMAT TMA
@@
 Transmission electron microscopy
-To examine the samples, high-resolution scanning transmission electron microscopy (HR-STEM) studies were carried out using a Themis Z(3.1) 80-300 TEM instrument (Thermo Fisher Scientific) with probe Cs-aberration correction, operating at 300 nm.
-The microscope is equipped with a Super-XTM (TFS) Energy Dispersive X-ray (EDX) detector and an a Gatan Image Filter (GIF) Continuum 1065ER used for Electron Energy Loss Spectroscopy (EELS).
-The samples were prepared by Focused Ion Milling (FIB) lift-out technique using a Helios G4 UC FIB-SEM device operated at 30, 16 and 5 kV.
+[…]
+Secondary Ion Mass Spectrometry
+[…]
 Ellipsometry
-The refractive indices were determined for each structure using an SE850 DUV variable angle spectroscopic ellipsometer (Sentech Instruments GmbH), and the data were evaluated in the wavelength range from 400nm400\ $\text${nm} to 1040nm1040\ $\text${nm} with a uniaxial Sellmeier dispersion model.
+The ordinary and extraordinary refractive indices were determined for each structure using an SE850 DUV variable angle spectroscopic ellipsometer (SENTECH Instruments GmbH), and the data were evaluated in the wavelength range from 400 to 1040nm1040\ $\text${nm} with a uniaxial Sellmeier dispersion model.
 Nonlinear setup
 The following optical setup was used to investigate the nonlinear optical response of the ABC layers.
-The femtosecond laser (PHAROS-SP, Light Conversion) emits pulses with full-width half-maxima (FWHM) of $177\text{\,}\mathrm{fs}$ with a central wavelength of 1032 nm and with an average power of up to $1.5\text{\,}\mathrm{W}$ at a repetition rate of $200\text{\,}\mathrm{kHz}$.
-Polarization-based measurements were performed by a femtosecond Satsuma laser (Amplitude) with full-width half maxima of $270\text{\,}\mathrm{fs}$, wavelength of 1031.4 nm and the repetition rate of $1\text{\,}\mathrm{MHz}$, Fig.
+A femtosecond laser (PHAROS-SP, Light Conversion) emits pulses with full-width half-maxima (FWHM) of $190(10)\text{\,}\mathrm{fs}$ with a central wavelength of 1032 nm and with an average power of up to $1.5\text{\,}\mathrm{W}$ at a repetition rate of $200\text{\,}\mathrm{kHz}$, see Fig.
 2.
-Power and polarization are controlled by a half-wave plate HW1 (WPH05M-1030, Thorlabs), polarizer P1 (VA5, Thorlabs), and half-wave plate HW2 (Thorlabs AHWP05M-950).
-The laser beam is focused on the sample by the plano-convex lens L1 (LA1172-B-ML, Thorlabs) with a focal length of 400 mm, producing a beam waist of $w_{0}=55\ \upmu\text{m}$.
-This leads to the confocal parameter of 18, that is, $2\timesRayleigh$ length.
-The loosely focused beam enables the front and back sides of the sample to remain in focus during the rotation process; the sample is 1 mm thick.
-A long pass filter LP (FELH0900, Thorlabs) is positioned behind the lens L1 to remove any parasitic light, especially any SHG created at other optical elements.
-The sample is mounted onto an automated rotation stage, and its position is precisely aligned so that the laser beam stays focused on the same position on the sample during the whole rotation.
-The beam spot must always lie on the y-axis of rotation.
-The sample is rotated from the angle of incidence $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $85\text{\,}\mathrm{\SIUnitSymbolDegree}$ in steps of $0.25\text{\,}\mathrm{\SIUnitSymbolDegree}$.
-This angle setting verifies proper sample alignment and improves the reliability of the evaluation process.
-To isolate the SH signal from the fundamental, a short-pass filter SP (FESH0800, Thorlabs) is used to remove the fundamental frequency.
-A Rochon prism analyzer P2 (RPM10, Thorlabs) is required to determine the direction of SHG polarization.
-A bandpass filter BP (Thorlabs FBH520-40) centered at 520 nm with FWHM 40 nm suppresses parasitic light and higher harmonics.
-The SHG signal is focused with a biconvex lens L2 (LA1608-A, Thorlabs) with a focal length of 100 mm on the camera (CS165MU1/M, Zelux).
-The camera calibration factor for calculating the SHG’s power is determined separately by using a Beta-Barium Borate (BBO) crystal as a reference sample.
-This samples produces a sufficiently strong SHG signal, which can be measured by a calibrated photodiode (S121C, Thorlabs) placed on the camera place.
-Subsequently, several optical density filters were introduced in the optical path to lower the power of the SH from the BBO crystal to the same order of magnitude as the one from the sample.
-Knowing the power output and optical density of the filters, we can correlate the power recorded by the camera with the measured signal.
-The signal was summed up in the area of images with recorded SHG, subsequently divided by the exposure time, and multiplied by the calibration factor of the camera to determine the optical power.
-Figure 2: Graphical schema of the optical setup used for second harmonic measurements of the ABC type heterostructures.
-From the left: femtosecond laser, HW1: half-wave plate, P1: polarizer, HW2: half-wave plate, L1: plano-convex lens, LP: long pass filter, rotation stage with the sample, SP: short pass filter, L2: biconvex lens, P2: Rochon prism analyzer, BP: bandpass filter, camera.
+The polarization dependence of the SHG was performed by another femtosecond laser (Satsuma, Amplitude) with a full-width half maxima of $270\text{\,}\mathrm{fs}$, wavelength of 1031.4 nm and repetition rate of $1\text{\,}\mathrm{MHz}$.
+Power and polarization were controlled using a half-wave plate HW1 (WPH05M-1030, Thorlabs), polarizer P1 (GT10-B, Thorlabs) and a half-wave plate HW2 (Thorlabs AHWP05M-950).
+The laser beam was focused on the sample using the plano-convex lens L1 (LA1172-B-ML, Thorlabs) with a focal length of 400 mm, producing a beam waist of $w_{0}=55(5)\ \upmu\text{m}$.
+This leads to a confocal parameter of 18 mm, that is, $2\timesRayleigh$ length.
+The loosely focused beam allows the front and back sides of the sample to remain in focus during tilting; the sample is 1 mm thick.
+A longpass filter LP (FELH0900, Thorlabs) was placed behind lens L1 to remove any parasitic light, especially any SHG created by the preceding optical elements.
+The sample was mounted on an automated rotation stage and aligned so that the laser beam remained focused on the same spot when the sample was tilted.
+The optical angle of incidence changed from $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $85\text{\,}\mathrm{\SIUnitSymbolDegree}$ in steps of $0.25\text{\,}\mathrm{\SIUnitSymbolDegree}$.
+The generated SH signal from the sample was separated from the fundamental using a shortpass filter SP (FESH0800, Thorlabs).
+For detecting the generated SH, the sample was imaged with a CMOS camera (CS165MU1/M, Thorlabs) using a biconvex lens with a focal length of 100 mm.
+Additionally, a band-pass filter (FBH520-40, Thorlabs) was used to suppress the environmental light.
+A Rochon prism analyzer P2 (RPM10, Thorlabs) was used to determine the direction of SHG polarization.
+The SH signal was focused using a biconvex lens L2 (LA1608-A, Thorlabs) with a focal length of 100 mm on the camera (CS165MU1/M, Zelux).
+The camera calibration factor to calculate the power of the SHG is determined separately using a beta-barium borate (BBO) crystal [3].
+Figure 2: Graphical schema of the optical setup used for second harmonic measurements of the ABC-type heterostructures.
+From the left: femtosecond laser, HW1: half-wave plate, P1: polarizer, HW2: half-wave plate, L1: plano-convex lens, LP: longpass filter, tilting stage with the sample, SP: shortpass filter, L2: biconvex lens, P2: Rochon prism analyzer, BP: bandpass filter, camera.
 Nonlinear characterization
-The standard method for evaluating nonlinear second-order susceptibility $\chi^{\textup{(2)}}$ is the Maker fringe technique Maker .
-It uses the rotation of the sample while the SHG intensity is collected for each angle.
+The standard method for evaluating nonlinear second-order susceptibility $\chi^{\textup{(2)}}$ is the Maker fringe technique [20].
+It uses the tilting of the sample and the SHG intensity is collected for each angle.
 However, it is limited to thicker crystals with a minimum size of tens of micrometers to a few millimeters.
-The limiting factor led to the development of a new method for measuring second-order susceptibility by Hermans et al.
-On-determination-SHG ; hermans2019 ; koskinen2018 .
-This method is well suited for evaluating the $\chi^{(2)}$ values of thin films with thicknesses even less than 100 nm where phase mismatch effects can be neglected.
-Like the Maker fringe technique, the sample is rotated during SHG measurements using a fixed polarization fundamental frequency, and the obtained SH curve is used to fit the nonlinear coefficients.
+This limiting factor led to the development of a new method for measuring second-order susceptibility by Hermans et al.
+[15, 14, 17].
+This method is well suited for evaluating the $\chi^{(2)}$ values of thin films with thicknesses even less than 100 nm, where phase mismatch effects can be neglected.
+Similarly to the Maker fringe technique, the sample was tilted during SHG measurements using a fixed polarization fundamental frequency, and the obtained SH curve was used to fit the nonlinear coefficients.
 This method does not consider individual separated layers A, B, and C and each interface as a source of second harmonic, but uses an effective medium approximation.
-Neglecting its birefringent properties, the ABC heterostructure is represented by an average refractive index of $(n_{0}+n_{e})/2$.
-The SHG is described in terms of the surface second-order nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ of the total heterostructure.
+Neglecting its birefringent properties, the ABC heterostructure was represented by an average refractive index of $(n_{\textup{o}}+n_{\textup{e}})/2$.
+SHG was described in terms of the surface second-order nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ of the total heterostructure.
 Any second-order nonlinear process is described by the second-order susceptibility tensor $\chi^{\textup{(2)}}$, with 27 elements.
@@
 The ABC heterostructure belongs to the symmetry group $C_{\infty v}$, as do other amorphous surfaces or thin achiral films with in-plane isotropy.
-This means that the nonzero coefficients of the second-order susceptibilities are $\chi^{\textup{(2)}}_{{xxz}}=\chi^{\textup{(2)}}_{{yyz}}=\chi^{\textup{(2)}}_{{zxz}}=\chi^{\textup{(2)}}_{{yzy}}$, $\chi^{\textup{(2)}}_{{zxx}}=\chi^{\textup{(2)}}_{{zyy}}$ and $\chi^{\textup{(2)}}_{{zzz}}$, where the $z$ direction is perpendicular to the layers and the $x$ and $y$ directions are longitudinal to the layers.
+This means that the nonzero coefficients of the second-order susceptibilities are $\chi^{\textup{(2)}}_{{xxz}}=\chi^{\textup{(2)}}_{{yyz}}=\chi^{\textup{(2)}}_{{zxz}}=\chi^{\textup{(2)}}_{{yzy}}$, $\chi^{\textup{(2)}}_{{zxx}}=\chi^{\textup{(2)}}_{{zyy}}$ and $\chi^{\textup{(2)}}_{{zzz}}$, where the $z$ direction is perpendicular to the layers and the $x$ and $y$ directions are parallel to the layers.
 It is also assumed that these coefficients have only real values and that the measurement takes place outside the resonance region, i.e.
 absorption is not present.
-The general description of the electric field components of the SH coming from a surface with $C_{\infty v}$ symmetry is described as:
+The general description [15] of the electric field components of the SH coming from a surface with $C_{\infty v}$ symmetry is described as:
 $E_{2\omega,\textup{p}}=fE_{\omega,\textup{p}}^{2}+gE_{\omega,\textup{s}}^{2},$ (1)
 $E_{2\omega,\textup{s}}=hE_{\omega,\textup{p}}E_{\omega,\textup{s}},$ (2)
-where p and s represent p- and s-polarized components and $E_{\omega}$ is the incoming electric field amplitude of the fundamental frequency.
+where p and s represent the p- and s-polarized components and $E_{\omega}$ is the incoming electric field amplitude of the fundamental frequency.
 The coefficients $f$, $g$, and $h$ represent generalized layer properties such as nonlinear susceptibility tensor components, thickness, refractive index, and angle of incidence.
@@
 We can obtain the bulk second-order susceptibility as $\chi^{\textup{ABC}}_{\textup{b}}=\chi^{\textup{ABC}}_{\textup{s}}/d$, where $d$ is the total thickness of the heterostructure.
-Under these approximations made by Hermans et al., a monochromatic p-polarized plane wave of the fundamental frequency of the amplitude of the electric field $E_{\omega,\textup{p}}$ produces a transmitted p-polarized SH electric field $E_{2\omega,\textup{total}}$ as
+Under these approximations made by Hermans et al.[15, 14, 17], a monochromatic p-polarized plane wave of the fundamental frequency of the amplitude of the electric field $E_{\omega,\textup{p}}$ produces a transmitted p-polarized SH electric field $E_{2\omega,\textup{total}}$ as
 $E_{2\omega,\textup{total}}$ $=E_{2\omega,\textup{front}}+E_{2\omega,\textup{back}}$ (3)
@@
 $\approx\left(\chi_{\textup{s},zxx}^{i}+2\chi_{\textup{s},xxz}^{i}\right)\sin(\Theta_{i})\cos^{2}(\theta_{i})+\chi_{\textup{s},zzz}^{i}\sin^{2}(\theta_{i})\sin(\Theta_{i}).$
-The parameters for the fundamental frequency (FF) $\omega$ are written in lower case letters, and for the SH in upper case.
+The parameters for the fundamental frequency (FF) $\omega$ are denoted by lower case letters, and for the SH by upper case.
 The parameters $T_{i,j}$ and $t_{i,j}$ are the Fresnel transmission coefficients for p-polarized light that transmits from medium $i$ to medium $j$.
-The variables $N_{i}$ and $n_{i}$ are refractive indices medium $i$.
+The variables $N_{i}$ and $n_{i}$ are refractive indices of medium $i$.
 The variables $\Theta_{i}$ and $\theta_{i}$ are angles of propagation in each medium $i$ with respect to the surface normal $z$.
@@
 The refractive indices for the glass substrate made of pure fused silica are $N_{\textup{glass}}=1.4615$ and $n_{\textup{glass}}=1.4500$.
-Approximation in Eq.
+The approximation in Eq.
 4 is valid only for materials with low dispersion with $\Theta_{i}\approx\theta_{i}$.
 Due to this approximation, the values $\chi_{\textup{s},zxx}^{\textup{ABC}}$ and $\chi_{\textup{s},xxz}^{\textup{ABC}}$ cannot be fitted independently, but only as a sum $A_{\textup{s},zx}^{\textup{ABC}}=\chi_{\textup{s},zxx}^{\textup{ABC}}+2\chi_{\textup{s},xxz}^{\textup{ABC}}$.
-For substrate nonlinearity we used literature values Kauranen_glass_calibration at wavelength 1064 nm $\chi_{\textup{s},zxx}^{\textup{glass}}=3.8\times 10^{-22}$ pm/V, $\chi_{\textup{s},xxz}^{\textup{glass}}=7.9\times 10^{-22}$ pm/V, $\chi_{\textup{s},zzz}^{\textup{glass}}=5.9\times 10^{-22}$ pm/V and we apply the Miller’s rule
+For substrate nonlinearity we used literature values [32] at the wavelength 1064 nm: $\chi_{\textup{s},zxx}^{\textup{glass}}=3.8\times 10^{-22}$ pm/V, $\chi_{\textup{s},xxz}^{\textup{glass}}=7.9\times 10^{-22}$ pm/V, $\chi_{\textup{s},zzz}^{\textup{glass}}=5.9\times 10^{-22}$ pm/V and we apply the Miller’s rule
 $\begin{aligned} \frac{\chi^{(2)}(2\omega)}{\chi^{(1)}(2\omega)\cdot\left|\chi^{(1)}(\omega)\right|^{2}}=\text{const.}\end{aligned},$ (5)
 with $\chi^{(1)}$ as the linear susceptibility of the material, to estimate the values at a wavelength of 1032 nm.
-However, since a femtosecond laser is used, a delay is introduced between the two pulses from the front and back sides of the sample.
-An SH pulse is generated on the front side of the sample.
-When it reaches the back side of the sample, another SH pulse is already generated because the pulse of the fundamental wavelength travels at different group velocity through the substrate compared to the SH pulse.
-This delay between the two SH pulses when they reach the backside of the sample is called the temporal walk-off time $t_{\text{walk-off}}$.
-This causes destructive interference to be incomplete, and thus, the depth of the interference fringes is decreased.
+Because a femtosecond laser is used, the group velocity mismatch between the fundamental and its SH introduces a temporal delay between the SH generated at the front and back surfaces of the sample.
+By the time the SH generated at the front surface reaches the back surface, an additional SH contribution has already formed there.
+This relative delay between the two SH contributions is known as the temporal walk-off time $t_{\text{walk-off}}$.
+This causes destructive interference to be incomplete, and thus, the contrast of the interference fringes is decreased.
 When the temporal walk-off is not taken into account, the value of $\chi_{\textup{}}^{(2)}$ is overestimated.
 For our fused silica sample, 1 mm thick, the value was measured to be $t_{\text{walk-off}}=94$ fs.
-This value corresponds well with previous measurements in the literature and also with the calculation of the delay based on the group velocity of each pulse in the substrate.
-Therefore, the model was modified by Hermans et al.
-for the pulsed laser with the temporal walk-off effect as
+This value corresponds also well with the calculation of the delay based on the group velocity of each pulse in the substrate.
+Therefore, the model was modified by Hermans et al.[15, 14, 17] for the pulsed laser with the temporal walk-off effect as
 $\begin{aligned} P_{2\omega}=K_{2}\int_{-\infty}^{+\infty}\left[E_{2\omega,\text{front}}\ \text{sech}^{2}\left(\frac{t}{\frac{\Delta t}{2\ln\left(1+\sqrt{2}\right)}}\right)+E_{2\omega,\text{back}}\ \text{sech}^{2}\left(\frac{t+t_{\text{walk-off}}}{\frac{\Delta t}{2\ln\left(1+\sqrt{2}\right)}}\right)\right]^{2}\,dt\end{aligned},$ (6)
@@
 Therefore, we keep the angle of incidence below $60^{\circ}$ for the analysis.
+For additional details, see Supplement 1.
 The primary source of uncertainty in the $\chi^{(2)}$ evaluation stems from the uncertainty in determining the angle of incidence, estimated as a systematic error of $\pm 1^{\circ}$.
@@
 The temporal walk-off effect also influences the interference by making the local minima shallower.
-III Results and discussion
-To investigate the nonlinear interface response of ABC heterostructures, multiple samples with a constant total deposition thickness of approximately 60 nm were fabricated, while systematically varying the ABC period thickness $t_{\text{ABC}}$ from 60 nm down to the atomic scale of 0.3 nm.
+3 Results and discussion
+Atomic layer deposition (ALD) enables precise control over thin-film thickness via self-limited growth, with growth rates per cycle (GPC) of $0.12$ nn/cycle for SiO2 (A), $0.07$ nn/cycle for TiO2 (B), and $0.15$ nn/cycle for Al2O3 (C).
+To investigate the nonlinear interface response of ABC heterostructures, multiple samples with a constant total deposition thickness of approximately 60 nm were fabricated, while systematically varying the ABC period thickness $t_{\text{ABC}}$ from 60 nm down to the atomic scale of 0.3 nm and it is easily achieved by ALD by controlling the number of cycles.
 This approach enabled the study of nonlinear responses by altering the density of interfaces per unit thickness and exploring different morphologies across the samples, with the goal of understanding how these variations influence the properties of heterostructures.
 Each oxide material is represented approximately equally in the period $t_{\text{ABC}}$.
-The ABC heterostructures consist of around 300–380 atomic layers for a 60 nm thick period and only about two atomic layers for an ABC period thickness of 0.3 nm, assuming a Si–O bond length of 0.15 nm, a Ti–O bond length of 0.20 nm, and an Al–O bond length of 0.16 nm Si-O ; Ti-O ; Al-O .
+The ABC heterostructures consist of around 300–380 atomic layers for a 60 nm thick period and only about two atomic layers for an ABC period thickness of 0.3 nm, assuming a Si–O bond length of 0.15 nm, a Ti–O bond length of 0.20 nm, and an Al–O bond length of 0.16 nm [24, 21, 38], suggesting ultimate physical limits for ABC layers.
 Scanning transmission electron microscope (STEM) images were collected using a high-angle annular dark-field (HAADF) and annular dark field (ADF) detectors to reveal the structural characteristics of the SiO2/TiO2/Al2O3 nanolaminates on a silicon wafer substrate in selected samples S3 and S2.
@@
 On the other hand, sample S2, with a reduced $t_{\text{ABC}}$ of 0.7 nm, shows no recognizable layer separation, indicating a loss of structural distinction on this scale.
-Figure 4: ABF STEM images showing the SiO2/TiO2/Al2O3 nanolaminates on the silicon wafer substrate of: (left) sample S3 with the ABC period thickness $t_{\textup{ABC}}$ of 1.5 nm; TiO2 layers appear darker; (right) sample S2 with $t_{\textup{ABC}}$ of 0.7 nm.
+Figure 4: ADF STEM images showing the SiO2/TiO2/Al2O3 nanolaminates on the silicon wafer substrate of: (left) sample S3 with the ABC period thickness $t_{\textup{ABC}}$ of 1.5 nm; TiO2 layers appear darker; (right) sample S2 with $t_{\textup{ABC}}$ of 0.7 nm.
 The layers are no longer visible.
@@
 5.
-The Bragg peaks are clearly visible for all samples except for the heterostructure with a period of 0.7 nm (sample S2).
+The Bragg peaks are clearly visible for all samples except for the heterostructure with a period of 0.7 nm (sample S2), suggesting notable layer intermixing and a lack of clearly separated layers.
 As the period of the heterostructure decreases, the first Bragg peak shifts to the higher grazing angle.
 The XRR analysis shows the existence of separated nanolaminates and the excellent agreement with the targeted ABC period thickness.
-The analysis also measured the total deposition thickness and provided the average interface width, incorporating both roughness and interlayer diffusion, see Table 2.
-Sample S2 has an interface width of 1.2 nm, greater than other samples, with a period thickness of $t_{\text{ABC}}=0.7$ nm, suggesting notable layer intermixing and a lack of clearly separated layers.
-Figure 5: Results of the X-ray reflectivity measurements proving the existence of the layer separation and verifying the period thickness $t_{\text{ABC}}$.
-Sample ID from the top: S13, S11, S6, S3, S2.
+The analysis also provided the total deposition thickness and provided the average interface roughness of all three types of interfaces in the heterostructure, see Table 2.
+Figure 5: Results of the X-ray reflectivity measurements proving the existence of the layers and verifying the period thickness $t_{\text{ABC}}$. Sample ID from the top: S13, S11, S6, S3, S2.
 Table 2: Results of the X-ray reflectivity measurements for selected SiO2/TiO2/Al2O3 heterostructures.
-Sample ID Total thickness $d$ (nm) Period thickness $t_{\textup{ABC}}$ (nm) Layer thickness SiO2 (nm) Layer thickness TiO2 (nm) Layer thickness Al2O3 (nm) Interface width (nm)
-S2 57.3 0.7 0.4 0.16 0.15 1.2
+Sample ID Total thickness $d$ (nm) Period thickness $t_{\textup{ABC}}$ (nm) Layer thickness SiO2 (nm) Layer thickness TiO2 (nm) Layer thickness Al2O3 (nm) Average interface roughness (nm)
 S3 59.7 1.5 0.4 0.4 0.7 0.4
@@
 S13 59.6 11.9 3.4 2.8 5.7 0.7
+Time-of-flight secondary ion mass spectrometry (TOF-SIMS) depth profiles of the heterostructures are shown in Fig.
+6.
+In sample S3 ($t_{\text{ABC}}=1.5$ nm), periodic modulation is observed only for Si+ ions, indicating distinct SiO2 layers, while the Al+ and Ti+ signals do not show a clear periodicity.
+In contrast, sample S11 ($t_{\text{ABC}}=5.9$ nm) exhibits periodic modulation of all constituent cations (Si+, Al+, Ti+).
+The lack of resolved Al2O3 and TiO2 layering in S3 does not rule out separated ABC layers, as it may arise from limited depth resolution due to differential sputtering yields, intermixing, or sample-preparation effects.
+Independent X-ray reflectivity (XRR) and scanning transmission electron microscopy (STEM) analyzes confirm the presence of separated layers in S3 with the 1.5 nm period.
+Figure 6: Depth profiles of positive secondary ion yields measured by TOF-SIMS in the ABC heterostructure: (a) sample S3 and (b) sample S11.
 Table 3: Sample parameters for SiO2/TiO2/Al2O3 heterostructures and their optical properties.
-Sample ID Deposition cycles SiO2/TiO2/Al2O3 Super cycles $M$ Layer thickness (nm) SiO2/TiO2/Al2O3 Period thickness $t_{\textup{ABC}}$ (nm) Total thickness $d$ (nm) $n_{\text{o}}$ (1032nm)(1032\ $\text${nm}) $n_{\text{e}}$ (1032nm)(1032\ $\text${nm}) $N_{\text{o}}$ (516nm)(516\ $\text${nm}) $N_{\text{e}}$ (516nm)(516\ $\text${nm}) $A_{\text{zx}}$ (pm/V) $\chi_{\text{zzz}}$ (pm/V)
-[…]
-[…]
-[…]
+Sample ID Deposition cycles SiO2/TiO2/Al2O3 Super cycles $M$ Layer thickness (nm) SiO2/TiO2/Al2O3 Period thickness $t_{\textup{ABC}}$ (nm) Total thickness $d$ (nm) $n_{\text{o}}$ (1032 nm) $n_{\text{e}}$ (1032 nm) $N_{\text{o}}$ (516 nm) $N_{\text{e}}$ (516 nm) $A_{\text{zx}}$ (pm/V) $\chi_{\text{zzz}}$ (pm/V)
+S1* 1:1:1 561 0.12/0.07/0.15 0.3 57 1.76 - 1.80 - 0.29 0.54
+S2* 2:3:2 80 0.24/0.21/0.3 0.7 57 1.81 - 1.86 - 0.50 0.96
+S3 4:6:4 40 0.48/0.42/0.6 1.5 59 1.77 1.75 1.83 1.81 0.80 2.0
 S4 4:6:4 120 0.48/0.42/0.6 1.5 174 1.79 1.75 1.84 1.80 0.68 2.1
 S5 4:6:4 160 0.48/0.42/0.6 1.5 233 1.78 1.76 1.84 1.80 0.77 2.5
-[…]
+S6 8:12:8 20 0.96/0.84/1.2 3.0 63 1.79 1.71 1.85 1.75 0.61 1.8
 S7 8:12:8 31 0.96/0.84/1.2 3.0 93 1.79 1.69 1.85 1.73 0.52 1.6
@@
 S10 8:12:8 70 0.96/0.84/1.2 3.0 215 1.78 1.71 1.84 1.75 0.69 1.9
+S11 16:24:16 10 1.92/1.68/2.4 6.0 60 1.77 1.66 1.83 1.71 0.15 0.86
+S12 16:24:16 20 1.92/1.68/2.4 6.0 114 1.79 1.64 1.84 1.67 0.070 0.96
+S13 32:48:32 5 3.84/3.36/4.8 12 60 1.78 1.64 1.84 1.68 0.090 0.70
+S14 80:120:80 2 9.6/8.4/12 30 59 1.82 1.68 1.88 1.71 0.092 0.52
 […]
-S12 16:24:16 20 1.92/1.68/2.4 6.0 114 1.79 1.64 1.84 1.67 0.070 0.96
-[…]
-[…]
-[…]
-*: No birefringence was measured, $\dagger$: The effective medium approximation is no longer valid in the wavelength range from 400 nm to 1040 nm based on the ellipsometric evaluation Layer and period thickness $t_{\textup{ABC}}$ is calculated based on the growth rate per cycle, and the total thickness $d$ is based on ellipsometric evaluation.
+*: No birefringence was observed.
+$\dagger$: The effective medium approximation is no longer valid in the wavelength range from 400 nm to 1040 nm based on the ellipsometric evaluation.
+Layer and period thickness $t_{\textup{ABC}}$ is calculated based on the growth rate per cycle, and the total thickness $d$ is based on ellipsometric evaluation.
 The uncertainty for the refractive index is $\pm 0.01$.
-Samples close to 60 nm of the total thickness are highlighted in gray.
-The dispersion curve of the ordinary and extraordinary refractive index is presented in Fig.
-6.
-The magnitude of birefringence, $\Delta n=n_{\text{e}}-n_{\text{o}}$, depends on the thickness of the ABC period, and completely disappears in samples with nonexistent layer separation.
+Samples with targeted 60 nm of the total thickness are highlighted in gray.
+ABC-type heterostructures have birefringent properties [14].
+For monocrystals, birefringence can be explained by the anisotropic electrical properties of the molecules that make up the crystals.
+However, birefringence can also be caused by anisotropy if the material is arranged on an order of magnitude larger than the size of the molecules and this distance is still smaller than the wavelength of light.
+The ABC-type heterostructure, composed of three materials arranged in thin layers, acts as an effective medium exhibiting form birefringence and functioning as a negative uniaxial crystal.
+At the same time, the entire structure is amorphous, and the optical axis is perpendicular to the layers [4, 9].
+Ellipsometry confirms the presence of the negative birefringence, and the dispersion curve of the ordinary and extraordinary refractive index are presented in Fig.
+7.
+The magnitude of birefringence, $\Delta n=n_{\text{e}}-n_{\text{o}}$, depends on the thickness of the ABC period, achieving up to $\Delta n=0.15$ for sample S12, and completely disappears in samples with nonexistent layers S1 and S2.
 Optical properties are summarized in Table 3.
-Figure 6: Ordinary and extraordinary refractive indices SiO2/TiO2/Al2O3 nanolaminates of the sample S3 show the presence of the form birefringence.
-Polarization measurements were carried out to verify the symmetry properties of the ABC stacks.
-Figure 7 shows s- and p-polarization of the SH signal from sample S3 as a function of the polarization angle of the fundamental field $\varphi$ with the angle of incidence $\vartheta=63^{\circ}$.
-S-polarization is represented by $\varphi=0^{\circ}$ and p-polarization by $\varphi=90^{\circ}$.
-The polarization measurements correspond with the symmetry group $C_{\infty v}$.
-The maximum SH signal is observed when both the fundamental and the second harmonic are p-polarized.
+Figure 7: Ordinary and extraordinary refractive indices of SiO2/TiO2/Al2O3 nanolaminates of the sample S3 show the presence of the form birefringence.
+Nonlinear polarization measurements were conducted to confirm the symmetry properties of ABC stacks.
+The linear polarization angle of the fundamental field, $\varphi$, was varied, while the second harmonic analyzer was fixed along the s- or p-polarization direction.
+At $\varphi=0^{\circ}$, the pump light is s-polarized, and at $\varphi=90^{\circ}$, it is p-polarized.
+Figure 8 displays the s- and p-polarized SH signals from sample S3 as a function of $\varphi$, with an incidence angle of $\vartheta=63^{\circ}$.
+The polarization data align with the $C_{\infty v}$ symmetry group, with the maximum SH signal observed when both the fundamental and SH fields are p-polarized.
+The discrepancy between the theoretical and experimental s-polarized SHG data may stem from the strain present in the ABC stack, which alters the nonlinear susceptibility tensor via the photoelastic effect [23].
 Figure 8: Normalized SHG power for s- and p-polarizations as a function of the polarization angle $\varphi$ of the fundamental corresponding with the symmetry group $C_{\infty v}$, measured for the sample S3.
+Dots represent measurements and lines normalized values according to the theory.
 Based on the experimental results for the p-polarized fundamental and the second harmonic in Fig.
-8(a), the intensity of the second harmonic increases with decreasing ABC period thickness from 60 nm until 1.5 nm, where a maximum is observed.
+9(a), the intensity of the second harmonic increases with decreasing ABC period thickness from 60 nm until 1.5 nm, where a maximum is observed.
 For periods thinner than 1.5 nm, the second harmonic signal decreases abruptly.
 As each layer is approximately 0.25 nm thick, intermixing at the interfaces likely degrades layer quality, and incomplete layer closure further reduces the SH signal.
-This hypothesis is supported by STEM imaging and XRR analysis of sample S2, which show no clear layer separation for an ABC period of $t_{\textup{ABC}}=0.7$ nm.
+This hypothesis is supported by STEM imaging and XRR analysis of sample S2, which show no clear layers for an ABC period of $t_{\textup{ABC}}=0.7$ nm.
 The nonlinear model for the determination of $\chi^{\textup{(2)}}$ is fitted to all SHG data across varying ABC periods, with selected curves shown in Fig.
-8(b).
-Figure 8: (a) Angle-dependent measurements of the second harmonic frequency from SiO2/TiO2/Al2O3 heterostructures deposited on 1 mm thick fused silica substrate.
+9(b).
+Figure 9: (a) Angle-dependent measurements of the second harmonic from SiO2/TiO2/Al2O3 heterostructures deposited on 1 mm thick fused silica substrate.
 The total deposited thickness is approximately 60 nm across all samples; only the morphology, i.e., the ABC period thickness $t_{\textup{ABC}}$, varies from sample to sample.
-Both FF and SH are p-polarized with average power $P=0.5$ W.
+Both $F$ with an average pump power $P=0.5$ W and SH are p-polarized.
 (b) Fitting of SHG data from selected samples to analyze the second-order susceptibility of the ABC heterostructures.
 The same trend is also reported for power dependency measurements.
-The samples are positioned at the angle of incidence when SHG maxima are observed for p-polarized FF and SHG.
-In addition, an uncoated substrate of fused silica with natural surface second-order nonlinearity is also included to compare it with the SH intensity across the ABC samples.
-Verification of second-harmonic generation is achieved by observing its quadratic power dependency, and excellent agreement is found by achieving a slope line of $2.026(9)$ for sample S3, Fig.
-9.
-The second harmonic exhibits a quadratic relationship with the ABC super cycles $M$, similar to the classical bulk crystals, Fig.
+The samples are aligned at the incidence angle that yields the maximum SHG signal for p-polarized fundamental light.
+For comparison, an uncoated substrate of fused silica with surface second-order nonlinearity is also included.
+Verification of second-harmonic generation is achieved by observing its quadratic power dependency, and excellent agreement is found by achieving a slope line of $2.026(9)$ for sample S3, see Fig.
 10.
-For each ABC period, second-order susceptibility is determined with the results shown in Fig.
-11(a).
+Several thicknesses were analyzed to clarify the properties at different deposited total thicknesses $d$.
+The second harmonic exhibits a quadratic relationship with the ABC super cycles $M$, similar to the classical bulk crystals, see Fig.
+11.
+Previously reported studies on ABC composites also identified the bulk-like SHG behavior [2, 7].
+For each ABC period, second-order susceptibility is determined by fitting with the results shown in Fig.
+12(a).
 The trend suggests an enhancement of both $\chi_{zzz}$ and $A_{zx}$ with decreasing ABC period thickness and achieving $\chi_{zzz}=2.0\pm 0.2$ pm/V for $t_{\textup{ABC}}=1.5$ nm.
 For shorter periods, the values of $\chi_{zzz}$ and $A_{zx}$ decrease rapidly.
-Several thicknesses were analyzed to clarify the properties at different deposited total thicknesses $d$.
-The analysis shows that the nonlinear susceptibility remains constant for total thicknesses from 60 nm to 233 nm for a period of 1.5 nm.
-A similar trend is also observed for periods of 3 nm and 6 nm.
-Figure 9: Second harmonic power dependence on fundamental power in SiO2/TiO2/Al2O3 heterostructures, measured at the angle of incidence of maximum second harmonic intensity for both p-polarized fundamental and second harmonic.
-Quadratic power dependence of the second harmonic for sample S3 with period $t_{\textup{ABC}}=1.5$ nm, fitted to the data.
-Figure 10: Second harmonic power scales with quadratic dependence on the film thickness and with the number of ABC super cycles $\text{ABC}]\times M$.
+The analysis also shows that the nonlinear susceptibility remains constant for total thicknesses from 60 nm to 233 nm for a period of 1.5 nm.
+A similar trend is also observed for periods of 3.0 nm and 6.0 nm in Fig.
+12(b).
+Figure 10: Average second harmonic power dependence on average fundamental power in SiO2/TiO2/Al2O3 heterostructures, measured at the angle of incidence of maximum second harmonic intensity for both p-polarized fundamental and second harmonic.
+Quadratic power dependence of the second harmonic for sample S3 with $t_{\textup{ABC}}=1.5$ nm, fitted to the data.
+Figure 11: Average second harmonic power scales with quadratic dependence on the film thickness and with the number of ABC super cycles $\text{ABC}]\times M$.
 Both fundamental and SH are p-polarized with average fundamental power $P=0.5$ W.
@@
 (b) Investigating second-order susceptibility $\chi_{zzz}$ across variations in ABC period and total film thickness.
-Our best ABC compositions are compared in Table 4 with other commonly used nonlinear crystals, as well as with previously reported results for ABC type heterostructures in the literature.
-We achieve higher nonlinearity than previously reported values, and we attribute this improvement to better-optimized PEALD processes.
+Our best ABC compositions are compared in Table 4 with other commonly used nonlinear crystals, as well as with previously reported results for ABC-type heterostructures in the literature.
+We achieve higher nonlinearity than previously reported values, and we attribute this improvement to higher quality thin film heterostructures, higher quality plasma source and ALD processes.
 For example, Alloatti et al.
-ABC-Alloatti report difficulties with the HfO2 process.
-Moreover, we achieved much thinner and still enclosed layers compared to the previous reports.
+[2] report difficulties with the HfO2 process and overall structural inhomogeneity of the heterostructure.
+Moreover, in this systematic study we achieved much thinner, smoother, and still enclosed layers compared to the previous reports [2, 7].
 ABC period thicknesses $t_{\textup{ABC}}$ were 2.1 nm and 2.7 nm, compared to our best sample with 1.5 nm.
@@
 Structure Material ABC material $n$ (-) $n_{\textup{o}}$$n_{\textup{e}}$ Point Group $\lambda$ (nm) $t_{\textup{ABC}}$ (nm) $d$ (pm/V) Ref.
-Monocrystal SiO2 (Quartz) - 1.53/1.54 $\textit{D}_{3}$ 1064 - $d_{\text{11}}$ = 0.3 lithiumbiobate ; quartz-n
-Monocrystal BaB2O4 (BBO) - 1.67/1.54 $\textit{C}_{3v}$ 1064 - $d_{\text{22}}$ = 2.3 lithiumbiobate ; BBO-n
-Monocrystal LiNbO3 (LN) - 2.23/2.16 $\textit{C}_{3v}$ 1064 - $d_{\text{33}}$ = -27 lithiumbiobate ; LN-n
-ABC Amorphous HfO2/TiO2/Al2O3 2.0/2.1/1.6 - $C_{\infty v}$ 800 2.7 $d_{\text{33}}$ = 0.4 On-determination-SHG
-ABC Amorphous In2O3/TiO2/Al2O3 2.2/2.1/1.6 - $C_{\infty v}$ 800 2.1 $d_{\text{33}}$ = 0.6 On-determination-SHG
+Monocrystal SiO2 (Quartz) - 1.53/1.54 $\textit{D}_{3}$ 1064 - $d_{\text{11}}$ = 0.3 [31, 10]
+Monocrystal BaB2O4 (BBO) - 1.67/1.54 $\textit{C}_{3v}$ 1064 - $d_{\text{22}}$ = 2.3 [31, 8]
+Monocrystal LiNbO3 (LN) - 2.23/2.16 $\textit{C}_{3v}$ 1064 - $d_{\text{33}}$ = -27 [31, 43]
+ABC Amorphous HfO2/TiO2/Al2O3 2.0/2.1/1.6 - $C_{\infty v}$ 800 2.7 $d_{\text{33}}$ = 0.4 [15]
+ABC Amorphous In2O3/TiO2/Al2O3 2.2/2.1/1.6 - $C_{\infty v}$ 800 2.1 $d_{\text{33}}$ = 0.6 [15]
 ABC Amorphous (S3) SiO2/TiO2/Al2O3 1.45/2.33/1.62 1.77/1.75 $C_{\infty v}$ 1032 1.5 $d_{\text{33}}$ = 1.0 This work
@@
 By convention, $\chi^{(2)}=2d$
-To further clarify the origin of the effective bulk nonlinearity, two identical ABC samples S3 were placed on top of each other.
-First, two samples with stacking sequence I) CBA…CBA-Silica/CBA…CBA-Silica and sequence II) CBA…CBA-Silica/Silica-ABC…ABC were investigated, Fig.
-12.
-Both stacking sequences produce almost the same maximum SHG intensity.
-However, the number of interference fringes is increased for sequence b), and the depth of fringes is reduced.
-The temporal walk-off effect is more significant because the optical thickness of the substrates between the heterostructures is twice as large.
-However, when the stacking sequence I) Silica-ABC…ABC/CBA…CBA-Silica is applied, the SHG almost vanishes.
-This result suggests that the centrosymmetry is restored.
+To further clarify the origin of the effective bulk nonlinearity [7], two identical ABC samples S3 were placed on top of each other.
+First, two samples with stacking sequence I) CBA…CBA-Silica/CBA…CBA-Silica and sequence II) CBA…CBA-Silica/Silica-ABC…ABC were investigated, see Fig.
+13.
+Both stacking sequences produce almost the same maximum SHG power.
+However, the number of interference fringes is increased for sequence II), and the contrast of the fringes is reduced.
+In this case, the temporal walk-off effect is more significant because the optical thickness of the substrates between the heterostructures is twice as large.
+However, when the stacking sequence III) Silica-ABC…ABC/CBA…CBA-Silica is applied, the SHG signal almost vanishes.
+This result suggests that the overall centrosymmetry of the stack is restored.
 However, SH waves must be generated in each ABC and CBA layer; therefore, they are generated in each stack with an opposite phase and destructively interfere.
-This leads to the conclusion that the two susceptibilities of the ABC and CBA layers must have an opposite sign, for example, $|\chi_{\textup{s}}^{\textup{ABC}}|=|-\chi_{\textup{s}}^{\textup{CBA}}|$.
-The polarity flipping of the second-order susceptibility by reversing the growing order was also demonstrated in semiconductor heterostructures based on quantum-engineered intersubband transitions, where the polarity is flipped by spatial inversion in the growth sequence chi2flipping , similar to our case.
+This leads to the conclusion that the two susceptibilities of the ABC and CBA layers must have an opposite sign, for example, $\chi_{\textup{s}}^{\textup{ABC}}=-\chi_{\textup{s}}^{\textup{CBA}}$.
+The polarity flipping of the second-order susceptibility by reversing the growing order was also demonstrated in semiconductor heterostructures based on quantum-engineered intersubband transitions, where the polarity is flipped by spatial inversion in the growth sequence [33], similar to our case.
 However, the ABC and CBA sequences can differ under real conditions.
@@
 Depending on the stacking sequence configuration, the SH response varies significantly.
-IV Conclusion
-We demonstrated an enhancement of the effective bulk second-order susceptibility in ABC type heterostructures by optimizing the thickness of the ABC period, achieving an optimal value of $t_{\text{ABC}}=1.5$ nm and a dominant tensor component of $\chi_{zzz}=2.0\pm 0.2$ pm/V, based on the second-order nonlinear response at the interfaces.
+Both $F$ with an average pump power $P=0.5$ W and SH are p-polarized.
+4 Conclusion
+We demonstrated an enhancement of the effective bulk second-order susceptibility in ABC-type heterostructures by optimizing the thickness of the ABC period, achieving a dominant tensor component of $\chi_{zzz}=2.0\pm 0.2$ pm/V at $t_{\text{ABC}}=1.5$ nm, based on the second-order nonlinear response at the interfaces.
 The observed non-linearity enhancement stems from the higher density of interfaces, symmetry breaking in the ABC stacks, and dielectric contrast between the materials.
-The nonlinearity could be further enhanced if the layers remain distinct and non-intermixed at sub-0.5 nm thicknesses, for example, by minimizing surface and interface roughness through deposition on ultra-smooth substrates.
-Experimental results indicate that the optimal strategy for maximizing second-order susceptibility is to incorporate the greatest number of ABC layers per unit thickness, while ensuring physical separation to maintain structural integrity.
+The nonlinearity could be further enhanced if the layers remain distinct and non-intermixed at the range of 0.5 nm, for example, by minimizing surface and interface roughness through deposition on ultra-smooth substrates.
+Extremely thin and homogeneous layers are needed.
+The evolution of the roughness inside the ALD multilayer stack is the result of both the replicated substrate roughness and the intrinsic film roughness.
+The coating material and process conditions can influence the latter.
+This may have the potential for further improvements, particularly in increasing layer density [34].
+Among other techniques, ALD has a high potential to provide such films for future advanced nonlinear photonic application.
+Experimental results indicate that the optimal strategy to maximize second-order susceptibility is to incorporate the greatest number of ABC layers per unit thickness, while ensuring physical separation to maintain structural integrity.
 However, the layer thickness is ultimately constrained by the limit of a single atomic layer.
-Therefore, future efforts should prioritize identifying a more suitable material composition, which remains challenging and elusive.
+Therefore, future efforts should also prioritize the identification of a more suitable material composition, which remains challenging and elusive.
 Additionally, ABC nanolaminates have potential for near-ultraviolet applications, leveraging their low absorption edge when constituent materials are carefully selected.
-Polarity flipping, combined with form birefringence, enables phase matching in bulk form-birefringent metamaterials formBirometamaterial .
-We envision that ABC type heterostructures will open new avenues for nonlinear optics by integration into nanophotonic waveguides to efficiently access the dominant tensor component $\chi_{zzz}$ ABC-waveguide , metasurfaces and other photonic platforms ABC-overcoating ; ABC-metasrufaceplasmonic .
-Funding
-This research received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the International Research Training Group (IRTG) 2675 “Meta-ACTIVE” (project number 437527638) and the Collaborative Research Center (CRC) 1375 “NOA.” Additional support was provided by the Fraunhofer Internal Programs under Grant No.
+Polarity flipping, combined with form birefringence, enables phase matching in bulk form-birefringent metamaterials [39].
+We envision that ABC-type heterostructures will open new avenues for nonlinear optics through integration into nanophotonic waveguides to efficiently access the dominant tensor component $\chi_{zzz}$ [12], and current efforts focus on integrating the heterostructures into nonlinear metasurfaces and other photonic platforms [16, 30].
+[…]
+Funding This research received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the International Research Training Group (IRTG) 2675 “Meta-ACTIVE” (project number 437527638) and the Collaborative Research Center (CRC) 1375, (project number 398816777) "NOA".
+Additional support was provided by the Fraunhofer Internal Programs under Grant No.
 SME 431 40-04871.
-We also gratefully acknowledge funding from the Federal Ministry of Research, Technology and Space under Project ID 13N16897 (Gradient).
-Authors also thank the European Regional Development Fund and the State of Brandenburg for the Themis Z microscope (part of Potsdam Imaging and Spectral Analysis (PISA) facility).
-Disclosures
-The authors declare no conflicts of interest.
-Data Availability
-Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
-Appendix A Appendix
-The following method On-determination-SHG was used to experimentally obtain the temporal walk-off time for our laser and 1 mm thick fused silica substrates.
-The femtosecond laser with p-polarization of the fundamental at a power of 1 W was used to measure the angle-dependent p-polarized SH signal from the front and back surfaces with known nonlinear susceptibility $\chi_{\textup{s}}^{\textup{glass}}$ on each side.
-Then, the equation is as follows:
-$E_{2\omega,\textup{total}}$ $=E_{2\omega,\textup{front}}+E_{2\omega,\textup{back}}$ (7)
-$=-j\frac{\omega}{2cN_{\textup{glass}}\cos(\Theta_{\textup{glass}})}t^{2}_{\textup{air,glass}}T_{\textup{glass,air}}E^{2}_{\omega,\textup{in}}$
-$\ \ \ \ \ \left(\chi_{\textup{s},xxz}^{\textup{glass}}\sin(2\theta_{\textup{glass}})\cos(\Theta_{\textup{glass}})+\chi_{\textup{s},zxx}^{\textup{glass}}\sin(\Theta_{\textup{glass}})\cos^{2}(\theta_{\textup{glass}})+\chi_{\textup{s},zzz}^{\textup{glass}}\sin^{2}(\theta_{\textup{glass}})\sin(\Theta_{\textup{glass}})\right)$
-$\ \ \ \ \Bigg[\exp\left(-j\frac{2\omega N_{\textup{glass}}\cos(\Theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)-\exp\left(-j\frac{2\omega n_{\textup{glass}}\cos(\theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)\Bigg].$
-After substituting into the Eq.
-6, we fit the parameters $K_{2}$, $L_{\textup{glass}}$ and $t_{\text{walk-off}}$.
-The fitting is restricted to the angle of incidence of 60° to limit the effect of multiple reflections in the sample.
-The obtained value for the temporal walk-off time is $t_{\text{walk-off}}=94$ fs, see Fig.
-13.
-Figure 13: The result of the fitting for the temporal walk-off time $t_{\text{walk-off}}$ for 1 mm fused silica substrate.
-Other parameters for the evaluation of the second-order susceptibility are shown in the Table 5.
-Table 5: Additional fitted parameters of SiO2/TiO2/Al2O3 heterostructures for the evaluation of the second-order susceptibility
-Sample ID Calibration constant $K_{2}$ (-) $\times 10^{3}$ $L_{\text{glass}}$ ($m$)
-S1 26 1.039
-S2 88 1.010
-S3 92 1.019
-S4 83 0.999
-S5 60 0.985
-S6 95 1.043
-S7 104 1.048
-S8 101 1.027
-S9 100 0.985
-S10 70 1.025
-S11 103 1.017
-S12 88 0.999
-S13 108 1.026
-S14 90 1.017
+We also gratefully acknowledge funding from the Federal Ministry of Research, Technology, and Space under Grant No.
+13N16897 (GRADIENT) and 13F1000A (PriFusio).
+The authors also thank the European Regional Development Fund and the State of Brandenburg for the Themis Z microscope (part of the Potsdam Imaging and Spectral Analysis (PISA) facility).
+[…]
+Disclosures The authors declare no conflict of interest.
+[…]
+Data Availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
+[…]
+Supplemental document See Supplement 1 for additional details of the nonlinear characterization.